### v2

# Investigating Feature and Model Importance in Android Malware Detection: An Implemented Survey and Experimental Comparison of ML-Based Methods

Abstract.
The popularity of Android means it is a common target for malware. Over the years, various studies have found that machine learning models can effectively discriminate malware from benign applications. However, as the operating system evolves, so does malware, bringing into question the findings of these previous studies, many of which report very high accuracies using small, outdated, and often imbalanced datasets. In this paper, we reimplement 18 representative past works and reevaluate them using a balanced, relevant, and up-to-date dataset comprising 124,000 applications. We also carry out new experiments designed to fill holes in existing knowledge, and use our findings to identify the most effective features and models to use for Android malware detection within a contemporary environment. We show that high detection accuracies (up to 96.8%) can be achieved using features extracted through static analysis alone, yielding a modest benefit ($\sim$1%) from using far more expensive dynamic analysis. API calls and opcodes are the most productive static and TCP network traffic provide the most predictive dynamic features. Random forests are generally the most effective model, outperforming more complex deep learning approaches. Whilst directly combining static and dynamic features is generally ineffective, ensembling models separately leads to performances comparable to the best models but using less brittle features.
Android, machine learning, static analysis, dynamic analysis, hybrid analysis
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Security and privacy Mobile platform security
1. Introduction
Smartphone usage has increased exponentially over recent years. According to Ericsson’s mobility report (Ericsson, 2023), there were 6.9 billion active smartphone subscriptions in 2023, with 70.48% of them using an Android smartphone (StatCounter, 2024). It is common practice to augment smartphones with applications, which in Android can be downloaded from the official Google Play Store (Play, 2024) or third-party application stores. The availability and easy access to applications through third party stores, in particular, provides attackers with a means of distributing malware. According to the latest report published by the computer security company G Data, a new piece of Android malware appears on the internet every 12 seconds (DATA, 2022).
Several approaches have been proposed to secure mobile operating systems (OS), including the application sandbox approach used by Android (Elenkov, 2015). In order to grant access to device services outside of the sandbox, Android uses a permissions system; however, this has its own shortcomings (Almomani and Khayer, 2020). A number of security solutions, including malware detectors, vulnerability detection, user and developer reviews have been proposed (Sufatrio et al., 2015). Of these solutions, Android malware detection, or anti-malware, is one of the most widely used. Anti-malware can initially be used to stop an application from being released into application stores or later at the user level, preventing the user from installing the application.
Traditional non-machine learning approaches to malware detection use signatures to detect malicious files. However, any slight variation to the file might cause the anti-malware to not detect the malware. In particular, this makes it difficult to detect zero-day attacks or existing malware variations. With the rapid growth of Android malware, there may be occasions when signature-based anti-malware will not be able to detect thousands, and potentially more, existing malware. Machine learning (ML) can be used to solve this problem. ML-based techniques require analysis of Android applications to extract features that represent the characteristics of the application. The analysis might be static or dynamic. Static analysis extracts features from the source code of the application, whereas dynamic analysis extracts features observed whilst monitoring the behaviour of the application in a running state (Sikorski and Honig, 2012).
We recently surveyed existing approaches that use machine learning to detect Android malware (Muzaffar et al., 2022). One of the insights from our study was that most of these approaches, including many of the recent ones, were evaluated using obsolete versions of Android running historic, and often quite small, collections of Android applications. This observation has also been made in a recent commentary piece (Guerra-Manzanares, 2023). This makes it difficult to determine both whether these approaches will work in the current Android ecosystem, and whether they will generalise to larger collections of Android applications. Since they were evaluated using a variety of different metrics on many different datasets, it is also difficult to infer meaningful comparisons regarding the importance of feature and model choices.
To address these problems, the aim of this study is to rigorously analyse the importance of feature and model choices while training ML models for Android malware detection. We primarily do this by reimplementing past works and reevaluating them using a large dataset that comprises current Android applications that we collected during the period of 2019–2021, allowing us to carry out a meaningful comparison of these methods within the current Android ecosystem and on a level playing-field. It is not feasible, nor perhaps desirable, to reimplement all the previous work in this area. Instead we focus on a group of studies which are representative of the diversity of modelling approaches used in this field, favouring those which have reported high levels of malware discrimination. This choice was guided by our previous review of published work on ML-based Android malware detection (Muzaffar et al., 2022). We also carry out a number of new experiments to fill in knowledge gaps in the existing literature, and consider the benefits of using ensemble models that combine existing approaches.
These are the main contributions of this work:
  • We review the Android feature space by reimplementing past works. To the best of our knowledge, this is the first-of-its-kind review and analysis of the importance of different static and dynamic features while training ML models.
  • We compile a balanced up-to-date dataset of Android malware and benign applications that is, at the time of writing this paper, the largest publicly available dataset for assessing Android anti-malware. We also share the tools  [^1] The dataset will be made publicly available once this paper is accepted. used to create this dataset to make it easier for the community to develop and assess future anti-malware.
  • Using this dataset, we reimplement and reevaluate past Android anti-malware approaches that used static, dynamic and hybrid feature analysis approaches.
  • We rigorously compare different types of ML models to determine which ones work best with particular Android feature sets.
  • We report the best performing features, models, and feature selection algorithms for Android malware detection.
  • We present an ensemble approach that leverages the best performing static and dynamic models, achieving an accuracy of 97.8% on our contemporary dataset.
The paper is organized as follows. Section 2 describes our methodology, including dataset collection, our static and dynamic analysis tools, and our evaluation framework. This methodology is then applied to static analysis, dynamic analysis, and hybrid analysis — the three main branches of feature analysis and model building in malware detection — in Sections 3, 4 and 5, respectively. Section 7 then considers ensemble methods that combine the best models from the previous sections. Section 8 discusses the main findings of the study, and Section 9 concludes.
2. Methodology
In this section, we present out framework for assessing Android anti-malware approaches. Section 2.1 describes the dataset we use, and how it was collected. Section 2.2 describes the tool we developed for performing static and dynamic analysis of Android applications. Section 2.3 outlines the metrics we use to assess ML models. Section 2.4 motivates the choice of features and our focus on binary classification. Section 2.5 presents a core set of ML models and feature selection methods which we use throughout the study.
2.1. Dataset Collection
The quality of the data used is a key factor in evaluating the quality of any ML model. In the context of assessing Android anti-malware approaches, we require a dataset that is both up-to-date and representative of the Android software ecosystem. We focus on the binary classification task of discriminating malware from benign software, and hence the dataset needs to contain samples of both malware and benign software. Ideally these two classes should be equal in number, since this mitigates against dealing with unbalanced data during the course of building and evaluating ML models.
We collected Android applications released from 2019-2021 from the stores and repositories which are typically used by end users. The lack of availability of scripts to build datasets prompted us to build crawlers to download these applications. Through periodic use, these platform-independent Python scripts can be used to maintain an up-to-date dataset. The stores we targeted were UpToDown (Uptodown, 2024), APKMirror (APKMirror, 2024) and F-Droid (F-Droid, 2024). The scripts crawl the websites for all applications irrespective of the categories the applications are filed under. We downloaded a total of 62,000 benign applications. To the best of our knowledge, this is the most realistic and up to date benign dataset available right now.
To collect malware, we used the most recent Android malware dataset from VirusShare (VirusShare.com, 2024) for malware released in 2020. We also identified malware while downloading applications from the application stores. We used VirusTotal (VirusTotal, 2024) reports to label all the applications. To prevent false positives in the benign dataset, we only included applications with zero positive tags from anti-malware in VirusTotal reports. The final malware dataset consisted of 62,000 applications, matching the size of the benign dataset.
2.2. DroidDissector
We developed an automated analysis tool, DroidDissector, to extract both static and dynamic features from Android applications. Drawing from the most prominent features highlighted in the literature, our tool attempts to provide coverage of the whole Android malware detection landscape (Muzaffar et al., 2023). The tool’s architecture is depicted in Figure 1. It comprises two sub-systems: one for extracting static features, and another for extracting dynamic features.
2.2.1. Static Analysis Sub-System
The static analysis sub-system takes in a directory of APK (the application file format used by Android) files as its parameter and produces static analysis reports for each of these applications.
APK files are first reverse-engineered and decompressed using APKtool (Connor Tumbleson, 2019) to decode the manifest file and smali files from the APK package. Permissions used by the application, the application package name and the different application components — activities, services, broadcast receivers, and content providers — are stated in the Android manifest file. APKtool decodes the dex (binary code) files into readable Java-like smali files. APKTool creates a directory for each APK file. The static analysis tool then analyzes the files produced by the APKTool to retrieve features. For each application, the features extracted by our static analysis tool include:
Hardware Components::
The hardware components that the application needs to use are defined in the manifest file of the application.
Requested Permissions::
The permissions required by the application as defined in the manifest file.
App Components::
The application components required by the application as defined in the manifest file. These are activities, services, broadcast receivers, and content providers.
Filtered Intents::
The filtered intents used by the application defined in the manifest file.
Used Permissions::
The permissions actually used by the application from manifest and smali files.
Network Addresses::
The network addresses present in the source code of the application from the smali files.
API Calls v30::
The API calls in the latest SDK present in the source code of the application in the smali files.
Restricted API Calls::
List of restricted API calls defined by Drebin (Arp et al., 2014). We built this feature set by analyzing the smali files extracted from the APK.
Suspicious API Calls::
List of suspicious API calls defined by Drebin (Arp et al., 2014). We built this feature set by analyzing the smali files extracted from the APK.
API Call Graphs::
FLOWDROID is used to extract call graphs produced by the application from smali files.
Opcodes::
The operation codes, or opcodes, present in the application. Opcodes are machine language instructions.
We built the static analysis tool using Python 3.8 and Java. Except for API call graphs, we performed the static analysis of each APK file using Python. We integrated FLOWDROID, a Java library for taint analysis, into our static analysis tool to extract API call graphs. A static analysis of an APK file can take anywhere from five seconds to ten minutes depending on the size of the application. We saved a separate file containing all the features extracted from an application for every APK file. The resulting feature files follow a standard format, comprising the feature type followed by “::” and then the name of the feature extracted. For example, “RequiredPermission::android.permission.ACCESS_NETWORK_STATE”. The tool stores API call graphs extracted using FLOWDROID in a separate JSON file. It also stores opcodes from an application’s bytecode in a separate file.
Figure 1. Automated analysis tool architecture
2.2.2. Dynamic Analysis Sub-System
The dynamic analysis sub-system runs on the readily-available Android emulators provided by Android SDK (Android Developers, 2024a). The virtual images in the emulators also allow root access, which is essential for dynamic analysis. Here we describe the environment required to run the tool and the features extracted by the tool.
2.2.3. Environment
We developed and tested the tool on Ubuntu, CentOS and Windows 10. The tool requires the following to run:
  • Python 3: All the dynamic modules require Python to run.
  • Frida-Android: The tool requires Frida’s Python library to hook to API calls during execution of an application.
  • Android Emulator: Our tool uses the emulator provided by Android SDK as the virtual environment to run the analysis. The emulator should be on the system path. We have tested the tool on Android 8, 9, 10 and 11 and we expect it to run on the upcoming Android 12.
  • The tool requires adb from Android SDK and it should be on the system path. The dynamic analysis tool connects to the virtual device using adb and executes commands on the emulator using adb.
  • Frida-server needs to be running on the virtual device to capture the hooked API calls while running the application.
  • The tool requires a static analysis report in order to analyse API calls. If a static analysis report is not provided, the dynamic analysis sub-system also requires APKTool to perform static analysis on the application.
2.2.4. Feature Extraction
The dynamic analysis sub-system can extract the key dynamic features used in Android malware detection. These include system calls, API calls, logcat files and network traffic. The tool uses Android’s monkey tool to input virtual events that imitate an application’s user input. It allows the user to set the number of virtual events and/or the maximum time the application should run for. The application is run in the controlled environment of the Android emulator provided by the Android SDK. There is an option to hide the emulator to save memory, but before running the analysis the Android emulator needs to be set up.
The dynamic analysis sub-system comprises the following feature extraction modules:
System Calls::
The system calls module extracts and saves all the system calls made by the application during its runtime using strace. This is a Linux utility used to monitor the interaction between the process (application) and the OS. While performing system calls analysis, it is essential to collect all the system calls that occur during runtime. From past works, we see researchers run the application and then retrieve the application process ID (pid) to monitor system calls associated with this. This approach might cause some system calls to be missed in the time between starting the application and running the strace tool. To mitigate against this, we run strace on the zygote process of the OS. The zygote process is the first process to run on an Android system. By monitoring the zygote process, we monitor all the processes in the OS. When the analysis ends, we filter out the system calls used by the application we are monitoring from the complete system call analysis extracted by monitoring the zygote process. The tool then saves the output in a separate file for every application.
Network Traffic::
Our tool uses Android emulator’s tcpdump to collect the complete network traffic of the application. It stores the resulting data in “pcap” files, later used to analyze and extract features. Wireshark (Wirshark, 2024) can process this file to analyze the application’s network traffic. It also saves a separate network traffic file for every application.
Android Logcat::
The tool uses logcat, a command line tool provided by Android SDK to dump all the messages and errors the device throws during the execution of the application. The adb utility runs logcat and saves the output in a file for further analysis.
API Calls::
The tool requires hooking of API calls beforehand to monitor execution of API calls during the run-time of an application. However, it is not computationally feasible to monitor the 130,000+ API calls present in the Android SDK. Therefore, the tool carries out the following steps: before starting the dynamic analysis, it checks if a corresponding static analysis report is present for the application. In addition to hooking the classic calls used in the literature of dynamic analysis tools (MobSF, 2023), we also hooked application-specific API calls. If present, it hooks the API calls extracted in the static analysis report, or else, it performs a static analysis on the application using the static analysis tool. The dynamic analysis tool then hooks the API calls to monitor during the execution of the application using Frida. The tool saves a separate file with all the APIs called during execution for further analysis. As the process may involve static analysis, the API call module is optional in the dynamic analysis tool. By default, the tool monitors a pre-defined list of API calls; however to monitor the API calls mined from an application’s static analysis, the static analysis module is required.
2.3. Evaluation Metrics
The following metrics are used to present results for the ML models trained in this study:
Confusion matrix:
The predictions of a malware detection model can be divided into four classes:
  1. (1):
    True Positive (TP): The application is correctly predicted as malicious.
  2. (2):
    False Positive (FP): The application is falsely predicted as malicious.
  3. (3):
    True Negative (TN): The application is correctly predicted as benign.
  4. (4):
    False Negative (FN): The application is falsely predicted as benign.
These four numeric values can be presented as a confusion matrix:
Class Positive Negative
Malware TP FN
Benign FP TN
Accuracy:
is the percentage of correct predictions.
(1) $Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$
Precision:
is the percentage of correctly classified malware from all predictions of malware.
(2) $Precision=\frac{TP}{TP+FP}$
True Positive Rate (TPR):
is the proportion of actual malware that are correctly identified. Also called recall.
(3) $TPR=\frac{TP}{TP+FN}$
F1-Score:
is the harmonic mean of precision and recall.
(4) $F_{1}=\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$
True Negative Rate (TNR):
is the proportion of actual benign applications that are correctly identified.
(5) $TNR=\frac{TN}{FP+TN}$
In general, it is necessary to use a range of metrics in order to give a full picture of the performance of different ML models (Lones, 2021). With this in mind, we report the accuracy, precision, F1-score, TPR and TNR for each model. TPR and TNR, in particular, help in understanding a model’s error rates. In the case of malware detection, it is particularly important to consider TPR, since a benign application labelled as malware is less likely to cause as many problems as malware labelled as benign.
The most common approach to comparing ML models in the Android malware literature is to report the average of metrics after using a resampling method like $k$-fold cross validation. Following this approach, we use standard 10-fold cross-validation (CV), and report the averages for all metrics. However, the average of a distribution can be misleading, particularly when there is a high variance between folds. To account for this, we report standard deviations and also carry out statistical tests when comparing averages of different approaches for malware detection. In a statistical test, a null hypothesis is set up (H0): the average is equal for all sample population groups. A significant difference between the samples is found if the results from the test were to reject the null hypothesis and accept the alternative hypothesis (Ha): the average is not equal for all sample population groups.
In this work, we use the Kruskal–Wallis (Kruskal and Wallis, 1952) omnibus test followed by pairwise Dunn’s (Dunn, 1964) tests. Kruskal–Wallis is a non-parametric test used to determine if there are significant differences between the medians of two or more distributions. In this case, the distributions are the set of $k$ metric scores produced by each ML model across $k$ folds. If the test statistic p-value is less than a threshold level, the null hypothesis is rejected, indicating that there is a significant difference between the groups. Kruskal-Wallis is an omnibus test and so does not indicate which sample group is statistically different from another. If Kruskal–Wallis finds a significant difference, Dunn’s test is then used to identify the samples that are statistically different from the others. If more than two models are being compared, a pairwise tournament between the ML models is used to determine the ML models with significant statistical differences. Dunn’s test automatically adjusts for multiple comparisons when used with Kruskal-Wallis. We use a threshold of p=0.05 for all statistical tests, i.e. a confidence of 95%.
2.4. Choice of Features and Classes
We used the tool described in Section 2.2 to extract features for Android malware detection. These features were chosen following a comprehensive review of the literature, as outlined in (Muzaffar et al., 2022), focusing on features which have been widely used in previous studies. We also took into account the findings of other prominent review papers (Naway and LI, 2018; Alqahtani et al., 2019; Bayazit et al., 2020; Wang et al., 2020).
In this paper, we focus on the binary classification task of discriminating malware from benign applications. However, some studies (e.g. Son et al. (2021, 2022)) consider multi-class problems, typically with the aim of discriminating between different malware families. We choose to focus on binary classification for several reasons. First, the aim of this study is to understand the role of different models and features in the malware detection process, and binary classification provides the most generalisable insights into this. Second, there is no universally-accepted method for labeling malware families, with the ever-changing and diverse nature of malware leading to varying taxonomies and classifications across the cybersecurity community. Again, this poses difficulties when trying to reach generalisable insights. Third, there is a significant imbalance within the natural distribution of malware types. For instance, in the dataset collected for this study, adware comprises around 70% of the samples. This introduces a challenge of handling imbalanced data, but also the challenge of separating the effects of mechanisms used to address imbalance from the effects of models and features.
2.5. Core ML Models and Feature Selection Algorithms
Previous studies have used a diverse range of ML models. In general, we reimplement the model(s) used in the original study wherever possible. However, to provide better consistency and comparability across studies, we also augment the original studies with the following core set of ML models, if they are not already included: support vector machine (SVM), random forest (RF), decision tree (DT) and $k$-nearest neighbour (kNN) (Hastie et al., 2009). All of these are widely used both within the Android malware detection literature (as surveyed in (Muzaffar et al., 2022)) and the broader ML literature. We use grid search to tune the hyperparameters of all the models implemented. This is to ensure fairness when making comparisons between models (Lones, 2021).
Many studies also use feature selection algorithms to reduce feature counts to a manageable level. Again, whenever possible, we reimplement the approach(es) used in the original study. However, we also augment these with the following core set of feature selection algorithms: mutual information (MI), chi-square, Pearson’s correlation coefficient (PCC), and variance threshold (VT). All of these are widely used in the literature.
3. Static Analysis
In this section, we focus on the use of static features in building Android malware detection models. Previous work in this area has considered various types of static features, including permissions, API calls, API call graphs and opcodes. We reimplement eleven of these previous studies, focusing on those that were notable either in terms of the methodologies they used or the performance of their models on the original datasets. We also report the results of new experiments that are intended to fill in some of the knowledge gaps found in previous studies.
3.1. Permissions and API calls
To begin with, we revisit one of the earliest studies that looked at permissions and API calls, carried out by Peiravian and Zhu (Peiravian and Zhu, 2013) in 2013. The authors compared models built using permissions alone, API calls alone, and a combination of both. In each case, they trained an SVM, a decision tree and a bagging classifier. They did not mention the kernel used for the SVM, nor the base classifier used for bagging, so in our reimplementation we used the default choices of a linear kernel for SVM and decision trees as the base classifier. At the time of their study, the total number of permissions in Android was 130 and the total number of API calls was only 1,326. As shown in Table 1, the numbers of both have grown since then, and by two orders of magnitude in the case of API calls.
Table 1. Comparison of samples and feature counts in Peiravian and Zhu’s and our dataset
Benign Malware Permissions API Calls
Peiravian and Zhu 1,250 1,260 130 1,326
Our dataset 62,000 62,000 166 134,207
Table 2. Result comparison between Peiravian and Zhu (Peiravian and Zhu, 2013) and our reimplementation, showing the effect of feature type (Perm: permissions alone, API: API calls alone, Combined: permissions and API calls) and ML model choice on performance metrics
Peiravian and Zhu. Results Our Results
Features (Classifier) Accuracy Precision F1-Score TPR TNR Accuracy Precision F1-Score TPR TNR
Perm (SVM) 0.935 0.924 - 0.875 - 0.935±$\pm 0.002$ 0.940±$\pm 0.002$ 0.933±$\pm 0.002$ 0.931±$\pm 0.002$ 0.941±$\pm 0.001$
API (SVM) 0.958 0.917 - 0.957 - 0.953±$\pm 0.004$ 0.957±$\pm 0.002$ 0.954±$\pm 0.001$ 0.956±$\pm 0.003$ 0.958±$\pm 0.001$
Combined (SVM) 0.969 0.957 - 0.948 - 0.956±$\pm 0.002$ 0.952±$\pm 0.002$ 0.954±$\pm 0.002$ 0.960±$\pm 0.001$ 0.952±$\pm 0.001$
Perm (DT) 0.924 0.898 - 0.866 - 0.901±$\pm 0.003$ 0.921±$\pm 0.003$ 0.918±$\pm 0.002$ 0.917±$\pm 0.004$ 0.921±$\pm 0.001$
API (DT) 0.933 0.894 - 0.903 - 0.945±$\pm 0.002$ 0.950±$\pm 0.002$ 0.954±$\pm 0.001$ 0.947±$\pm 0.002$ 0.950±$\pm 0.002$
Combined (DT) 0.945 0.906 - 0.928 - 0.949±$\pm 0.002$ 0.949±$\pm 0.003$ 0.949±$\pm 0.001$ 0.950±$\pm 0.001$ 0.949±$\pm 0.002$
Perm (Bagging) 0.936 0.920 - 0.882 - 0.930±$\pm 0.002$ 0.932±$\pm 0.001$ 0.933±$\pm 0.003$ 0.931±$\pm 0.001$ 0.932±$\pm 0.002$
API (Bagging) 0.949 0.936 - 0.907 - 0.961±$\pm 0.002$ 0.960±$\pm 0.003$ 0.959±$\pm 0.002$ 0.961±$\pm 0.002$ 0.960±$\pm 0.002$
Combined (Bagging) 0.964 0.949 - 0.941 - 0.948±$\pm 0.001$ 0.949±$\pm 0.002$ 0.945±$\pm 0.002$ 0.946±$\pm 0.001$ 0.948±$\pm 0.002$
Table 2 reports the results both for Peiravian and Zhu’s original study and for our reimplementation using our much larger contemporary dataset. First of all, it is interesting to see that, despite the much larger feature and data set sizes, our results cover a similar range of accuracies. Peiravian and Zhu provided figures for accuracy, precision and recall in their evaluation. For completeness, we report the full set of metrics for our reimplementation.
Peiravian and Zhu found that models based only on API calls outperformed models based only on permissions. The figures in Table 2 show that this also appears to be the case for our reimplementation, with an even larger margin than in the original study.
We analysed the differences between models further by applying statistical tests. First, Kruskal-Wallis showed that there are significant differences in the accuracies reported in Table 2 (H = 50.04, p $<$ 0.05). Post hoc Dunn’s tests then showed that (i) the bagging and DT models using API calls perform better, on average, than all models that use permissions, and (ii) SVM models using API calls perform better, on average, than SVM models that use permissions. This supports the conclusion that API calls are more useful than permissions for building Android malware detection models.
All other group-wise differences were not significant at the 95% confidence level, suggesting that the exact choice of ML model is less important. However, we note that, for API calls, the bagging model had the best performance across all metrics, achieving an average accuracy of 96%.
3.2. A Closer Look at Permissions
Nevertheless, permissions are easier to extract than other static features. For this reason, they are widely used in practice. To provide more insight into the best way to use permissions for Android malware detection, we reimplemented a study by Wang et al. (Wang et al., 2014) that aimed to provide a more fine-grained view of the risk associated with permissions. Their study used two groups of Android permissions to build ML models: normal permissions, which are automatically granted by the system, and dangerous permissions, which provide access to private data or allow control over the device. For more information about these categories, see (Android Developers, 2024b). As shown in Table 3, the total number of normal and dangerous permissions is only slightly different in the current Android release.
Their study used four feature selection algorithms to identify the permissions that are most useful for building ML models: MI, PCC, sequential forward selection (SFS), and T-tests. They also used principal component analysis (PCA) as an alternative means of dimensionality reduction. In our reimplementation, we trained our four base ML models using feature sets derived both from the feature reduction methods used by Wang et al. and also our standard set of feature selection methods. We generated feature sets of sizes 10 to 70 features using each method.
Table 3. Comparison of samples and total number of normal and dangerous permissions in Wang et al’s and our dataset
Benign Malware Permissions
Wang et al. 310,926 4,868 88
Our dataset 62,000 62,000 82
Table 4. Result comparison between Wang et al. (Wang et al., 2014) and our reimplementation, showing the effect of feature selection algorithm choice. The original study used SVM; for the results of our reimplementation, we also show the best-performing model in each case. Accuracy and F1-scores for the original study are approximate values as they were read from plots provided by the authors.
Wang et al. Results Our Results
Method Accuracy Precision F1-Score TPR TNR Method Accuracy Precision F1-Score TPR TNR
MI 0.996 - 0.895 0.923 - MI (RF) 0.930±$\pm 0.001$ 0.932±$\pm 0.001$ 0.929±$\pm 0.002$ 0.931±$\pm 0.002$ 0.931±$\pm 0.003$
PCC 0.996 - 0.895 0.923 - PCC (RF) 0.907±$\pm 0.002$ 0.910±$\pm 0.003$ 0.905±$\pm 0.002$ 0.902±$\pm 0.003$ 0.911±$\pm 0.001$
T-Test 0.996 - 0.895 0.923 - T-Test (RF) 0.919±$\pm 0.001$ 0.920±$\pm 0.002$ 0.921±$\pm 0.002$ 0.919±$\pm 0.002$ 0.921±$\pm 0.001$
SFS 0.996 - 0.895 0.923 - SFS (SVM) 0.902±$\pm 0.002$ 0.902±$\pm 0.003$ 0.901±$\pm 0.002$ 0.901±$\pm 0.001$ 0.902±$\pm 0.002$
PCA 0.996 - 0.895 0.926 - PCA (SVM) 0.928±$\pm 0.001$ 0.926±$\pm 0.002$ 0.927±$\pm 0.003$ 0.932±$\pm 0.001$ 0.925±$\pm 0.002$
Chi-square - - - - - Chi-square(RF) 0.927±$\pm 0.002$ 0.923±$\pm 0.003$ 0.946±$\pm 0.002$ 0.901±$\pm 0.003$ 0.949±$\pm 0.003$
Table 4 shows the best results for each of Wang et al’s feature reduction methods, and for chi-square, which was our overall best performing feature selection method in terms of F1-score. The results from Wang et al. are for SVM models. For our results, we also indicate the best performing ML model for each feature reduction method. Our models performed significantly better than Wang et al’s in terms of F1-score. Although Wang et al. reported very high accuracies in their study, this is likely to be a reflection of their very imbalanced data set, and possibly an indication that their models overfit the majority class. This would also explain why their F1-scores were relatively low.
For our models, a positive Kruskal-Wallis test (H = 30.62, p $<$ 0.05) followed by pairwise Dunn’s tests showed that the only significant difference in mean accuracies was between the SVM trained on the SFS feature set and the other models. This does not allow us to draw any firm conclusions about which model or feature selection method is best.
Figure 2. Relationship between number of permissions used and model accuracy when permissions are ranked using different methods
Fig. 2 shows how the accuracy of each model changes based on the size of the feature set. It can be seen that, for all models and feature selection methods, accuracy initially improves as the number of features is increased, but then reaches a plateau point. The best overall accuracy was achieved by RF and SVM models when using 40 permissions, and using more than 40 permissions did not lead to improvement in most cases. A notable difference from the findings of Wang et al. is that this plateau point comes later, which may indicate that more permissions are now required to build reliable malware detection models. Interestingly, the lists of the top 40 permissions produced by MI, PCC and T-test had 38 permissions in common across all the CV test folds. For chi-square, the best performing feature set, of 60 permissions, included all the permissions chosen by these other feature selection algorithms. Moreover, the models trained using chi-square permissions plateau later than the other feature selection models and so require more permissions to achieve high accuracy. The best chi-square model also achieved the highest F1-score.
Rathore et al. (Rathore et al., 2021) further experimented with different feature selection and feature dimensionality reduction methods, with the aim of building an efficient model that used a minimal set of permissions. Apart from the original dataset with full permissions (OD), they used VT, PCA, and autoencoders (with one and three layers: AE-1L,AE-3L) to construct reduced feature sets. To this, we added our core set of feature selection algorithms. Unlike Wang et al., they carried out feature selection on the entire set of Android permissions, not just the normal and dangerous permissions. They trained six ML models for each of the feature selection/reduction methods: DT, kNN, SVM, RF, AdaBoost and DNNs (with two, four and seven layers). These were compared against models trained on the full set of 166 permissions.
Table 5. Result comparison between Rathore et al. (Rathore et al., 2021) and our reimplementation, showing the effect of model choice for permissions-based models, and the most effective feature selection method for reducing the number of permissions for each ML model type
Rathore et al. Results Our Results
Classifier Method Accuracy Precision F1-Score TPR TNR Method Accuracy Precision F1-Score TPR TNR
Decision Trees OD 0.926 - - 0.919 - OD 0.901±$\pm 0.003$ 0.921±$\pm 0.003$ 0.918±$\pm 0.002$ 0.917±$\pm 0.004$ 0.921±$\pm 0.001$
kNN PCA 0.914 - - 0.913 - MI 0.932±$\pm 0.002$ 0.928±$\pm 0.001$ 0.931±$\pm 0.002$ 0.935±$\pm 0.001$ 0.928±$\pm 0.002$
SVM OD 0.910 - - 0.894 - AE-3L 0.933±$\pm 0.002$ 0.927±$\pm 0.002$ 0.935±$\pm 0.002$ 0.942±$\pm 0.003$ 0.927±$\pm 0.001$
Random Forest OD 0.940 - - 0.930 - VT 0.949±$\pm 0.003$ 0.948±$\pm 0.002$ 0.950±$\pm 0.001$ 0.951±$\pm 0.001$ 0.949±$\pm 0.002$
AdaBoost AE-1L 0.911 - - 0.898 - OD 0.928±$\pm 0.003$ 0.922±$\pm 0.001$ 0.927±$\pm 0.001$ 0.934±$\pm 0.001$ 0.921±$\pm 0.002$
DNN-2L AE-1L 0.931 - - 0.914 - MI 0.930±$\pm 0.002$ 0.922±$\pm 0.003$ 0.933±$\pm 0.002$ 0.938±$\pm 0.003$ 0.921±$\pm 0.001$
DNN-4L VT 0.931 - - 0.920 - PCA 0.932±$\pm 0.001$ 0.923±$\pm 0.002$ 0.931±$\pm 0.001$ 0.941±$\pm 0.003$ 0.922±$\pm 0.002$
DNN-7L OD 0.930 - - 0.928 - OD 0.927±$\pm 0.002$ 0.922±$\pm 0.001$ 0.925±$\pm 0.002$ 0.933±$\pm 0.003$ 0.921±$\pm 0.002$
Figure 3. Effect of changing the variance threshold upon accuracy of random forest models. Also shows the number of permissions selected for different variance thresholds.
Table 5 summarises the performance metrics from our reimplementation, showing the figures for the most effective feature reduction method for each ML model. Our results are generally quite similar to those from the original study. Like Rathore et al., we found RFs to be the most effective model, with the highest mean values across all metrics. This is supported by a statistical analysis of the results in Table 5, where Dunn’s tests following a positive Kruskal-Wallis test (H = 54.58, p $<$ 0.05) indicated that RF accuracy was significantly higher than all other models except kNN. Our results also show that there is no particular value in reducing the set of permissions to those which are normal and dangerous prior to carrying out feature selection.
We found that RFs work best when combined with variance threshold. Digging down a bit further, Fig. 3 shows the effect of changing the threshold, indicating that lower thresholds work better. For a threshold of 0.2, the number of permissions can be reduced by 64%, whilst still producing better malware detection than when using all permissions.
We reimplemented one more study that used only permissions. This study, by Sahin et al. (Şahin et al., 2023), is notable for using model-based feature selection. Unlike the more traditional forms of feature selection used in the previous studies, model-based feature selection uses an ML model to determine the importance of each feature. This can lead to more appropriate feature combinations; however, it does have a cost, and Sahin et al. attempt to mitigate against this by using a relatively inexpensive model, namely multiple regression, to assess features.
Table 6. Result comparison between Sahin et al. (Şahin et al., 2023) and our reimplementation, showing the effect of model choice when permissions-based features are selected using a model-based feature selection algorithm
Sahin et al. Results Our Results
Model Permissions Accuracy Precision F1-Score TPR TNR Permissions Accuracy Precision F1-Score TPR TNR
SVM - - - - - - 60 0.914±$\pm 0.001$ 0.912±$\pm 0.001$ 0.913±$\pm 0.003$ 0.921±$\pm 0.003$ 0.911±$\pm 0.003$
Random Forest 27 - - 0.930 - - 80 0.924±$\pm 0.003$ 0.920±$\pm 0.003$ 0.925±$\pm 0.003$ 0.932±$\pm 0.003$ 0.921±$\pm 0.001$
Decision Trees 27 - - 0.956 - - 80 0.909±$\pm 0.001$ 0.914±$\pm 0.001$ 0.911±$\pm 0.001$ 0.909±$\pm 0.003$ 0.911±$\pm 0.001$
kNN 43 - - 0.954 - - 60 0.912±$\pm 0.003$ 0.912±$\pm 0.003$ 0.907±$\pm 0.003$ 0.911±$\pm 0.001$ 0.922±$\pm 0.003$
Naïve Bayes 43 - - 0.928 - - 20 0.817±$\pm 0.003$ 0.824±$\pm 0.002$ 0.866±$\pm 0.002$ 0.913±$\pm 0.001$ 0.905±$\pm 0.003$
Table 6 summarises the results of our reimplementation. These show that using a model-based feature selection approach based around multiple regression does not lead to an improvement in classification performance. This conclusion is backed by a statistical analysis (Kruskal-Wallis test (H = 31.22, p $<$ 0.05) of the results in Table 6 and the findings from the top-performing permission-based models listed in Table 1. It is possible that better performance could be obtained by matching the model used for classification with the model used for feature selection; however, this benefit was not observed in the original study, and the use of more expensive models for feature selection would create challenges for the larger feature sets considered in the remainder of this study.
To conclude, reasonable levels of performance can be achieved with permissions-based approaches, especially when the number of permissions is reduced using a feature selection algorithm, and RFs are used as the ML model. However, none of the permissions-based models we looked at were competitive against API call-based models.
3.3. Representations of API Calls
Next, we take a closer look at the best way of building models based on API calls, beginning with a reimplementation of the research carried out by Ma et al. (Ma et al., 2019), who considered three different ways of representing feature usage within an application: API usage, API frequency and API sequences.
For API usage, each API call is represented by a binary feature indicating whether the call is used within an application. We created our version of this dataset by using the complete set of 134,207 API calls; i.e., each application is represented by a feature vector $A=\{API_{1},API_{2},API_{3},\dots API_{134207}\}$, making this a high-dimensional feature set. This is the same dataset used for our earlier reimplementation of Peiravian et al. (Peiravian and Zhu, 2013). The API frequency data set is similar, except that each feature has an integer value, and indicates how many times each call is made within an application. For example, the feature vector for an application would be $M=\{5,4,0\}$ if $A=\{API_{1},API_{2},API_{3}\}$ and the application used $API_{1}$ 5 times, $API_{2}$ 4 times and $API_{3}$ 0 times.
For API sequences, each application is represented as a variable-length sequence of API calls. In our reimplementation, a call graph is extracted using the FLOWDROID module of our static analysis tool. The tool stores the APIs in chronological order and hence allows us to gather information on the order in which the application calls the APIs. We developed an algorithm which uses depth-first search (DFS) to obtain a set of routes for an application. The last API of every route is usually a system API call as user methods and user-defined APIs will always call system APIs. Therefore, we were able to extract system APIs that the application called in chronological order making an API sequence dataset. The resulting API system calls sequence is then converted into a numerical representation. For example, if we have a set of APIs $A=\{a,b,c,d\}$ then the applications are subsets of $A$. If $M=[a,c,a,d]$ is a malware application and $B=[d,c,b,a]$ is a benign application, then the vector of $M$ will be $\{1,3,1,4\}$ and the vector of $B$ will be $\{4,3,2,4\}$, where 1 represents $a$, 2 represents $b$, etc.
Table 7. Result comparison between Ma et al. (Ma et al., 2019) and our reimplementation, comparing the four different API call modelling approaches used in the original study
Ma et al. Results Our Results
Dataset Model Accuracy Precision F1-Score TPR TNR Model Accuracy Precision F1-Score TPR TNR
Usage DT - 0.968 0.965 0.962 - Random Forest 0.966±$\pm 0.002$ 0.968±$\pm 0.002$ 0.966±$\pm 0.002$ 0.964±$\pm 0.003$ 0.969±$\pm 0.001$
Frequency DNN(8) - 0.977 0.974 0.971 - DNN (16) 0.970±$\pm 0.001$ 0.971±$\pm 0.001$ 0.969±$\pm 0.002$ 0.979±$\pm 0.002$ 0.971±$\pm 0.002$
Sequence LSTM(8) - 0.985 0.986 0.988 - LSTM (8) 0.935±$\pm 0.003$ 0.935±$\pm 0.002$ 0.936±$\pm 0.001$ 0.936±$\pm 0.002$ 0.935±$\pm 0.001$
Ensemble Voting - 0.991 0.99 0.988 - Voting 0.965±$\pm 0.002$ 0.963±$\pm 0.003$ 0.962±$\pm 0.003$ 0.967±$\pm 0.002$ 0.963±$\pm 0.002$
Table 8. The effect of varying the number of layers in the deep neural network for DNN models trained on API usage, API frequency and API sequence feature sets
API Dataset Layers Accuracy Precision F1-Score TPR TNR
Usage 4 (DNN) 0.959±$\pm 0.002$ 0.958±$\pm 0.002$ 0.958±$\pm 0.002$ 0.957±$\pm 0.003$ 0.954±$\pm 0.002$
Usage 8 (DNN) 0.953±$\pm 0.002$ 0.952±$\pm 0.003$ 0.954±$\pm 0.002$ 0.955±$\pm 0.003$ 0.954±$\pm 0.002$
Usage 16 (DNN) 0.961±$\pm 0.001$ 0.962±$\pm 0.003$ 0.963±$\pm 0.002$ 0.959±$\pm 0.001$ 0.962±$\pm 0.001$
Usage 32 (DNN) 0.960±$\pm 0.001$ 0.959±$\pm 0.002$ 0.961±$\pm 0.002$ 0.958±$\pm 0.003$ 0.961±$\pm 0.001$
Frequency 4 (DNN) 0.961±$\pm 0.001$ 0.955±$\pm 0.002$ 0.958±$\pm 0.002$ 0.963±$\pm 0.002$ 0.956±$\pm 0.002$
Frequency 8 (DNN) 0.967±$\pm 0.002$ 0.968±$\pm 0.003$ 0.966±$\pm 0.002$ 0.966±$\pm 0.002$ 0.968±$\pm 0.002$
Frequency 16 (DNN) 0.970±$\pm 0.001$ 0.971±$\pm 0.002$ 0.969±$\pm 0.002$ 0.969±$\pm 0.002$ 0.971±$\pm 0.002$
Frequency 32 (DNN) 0.960±$\pm 0.002$ 0.961±$\pm 0.002$ 0.961±$\pm 0.002$ 0.962±$\pm 0.002$ 0.962±$\pm 0.002$
Sequence 4 (LSTM) 0.888±$\pm 0.002$ 0.890±$\pm 0.002$ 0.884±$\pm 0.002$ 0.881±$\pm 0.001$ 0.891±$\pm 0.003$
Sequence 8 (LSTM) 0.936±$\pm 0.003$ 0.935±$\pm 0.004$ 0.934±$\pm 0.001$ 0.936±$\pm 0.002$ 0.935±$\pm 0.001$
Sequence 16 (LSTM) 0.929±$\pm 0.001$ 0.930±$\pm 0.002$ 0.926±$\pm 0.002$ 0.927±$\pm 0.002$ 0.931±$\pm 0.002$
Sequence 32 (LSTM) 0.927±$\pm 0.002$ 0.927±$\pm 0.001$ 0.928±$\pm 0.002$ 0.929±$\pm 0.002$ 0.927±$\pm 0.002$
Table 9. Comparison of different ML models using the API frequency feature set
Classifier Accuracy Precision F1-Score TPR TNR
SVM 0.951±$\pm 0.001$ 0.952±$\pm 0.002$ 0.953±$\pm 0.002$ 0.949±$\pm 0.002$ 0.951±$\pm 0.002$
Random Forest 0.962±$\pm 0.001$ 0.963±$\pm 0.002$ 0.961±$\pm 0.002$ 0.958±$\pm 0.001$ 0.962±$\pm 0.003$
DT 0.942±$\pm 0.001$ 0.943±$\pm 0.002$ 0.941±$\pm 0.002$ 0.944±$\pm 0.001$ 0.945±$\pm 0.001$
kNN 0.943±$\pm 0.001$ 0.944±$\pm 0.002$ 0.945±$\pm 0.003$ 0.948±$\pm 0.001$ 0.949±$\pm 0.003$
DNN 0.970±$\pm 0.001$ 0.971±$\pm 0.001$ 0.969±$\pm 0.002$ 0.979±$\pm 0.002$ 0.971±$\pm 0.002$
Ma et al. (Ma et al., 2019) used a different model for the three datasets: a DT for the API usage dataset, a deep neural network (DNN) for the API frequency dataset, and a long short-term memory (LSTM) model for the API sequence dataset. They also built an ensemble from the three models. In Table 7, we summarise our results from using this same approach. This indicates that the API frequency model performs the best, with both the usage and frequency models outperforming the API sequence models. We did not find any benefit to ensembling the models. This is in opposition to the findings of Ma et al. (Ma et al., 2019), for whom the best model was the ensemble, followed by the sequence model. Like Ma et al., we also looked more closely at the effect of network depth on the performance of the DNN and LSTM models. Table 8 shows the results, showing that best performance is now achieved with a DNN of 16 layers, which is twice the depth of the best DNN found by Ma et al.
Kruskal-Wallis showed that the differences between the average accuracies in Table 7 are significant (H = 35.63, p $<$ 0.05). Dunn’s tests indicated a significant difference between the frequency model and the sequence model, and no significant difference between the ensemble model, usage model and the frequency model. This supports the conclusion that the simpler usage and frequency-based models are more discriminative, at least within a contemporary Android environment.
A limitation of the original study was that different models were used for different representations, meaning that the effect of changing the representation could not be isolated from the effect of changing the model. To give more insight into whether the additional information embedded in the frequency representation over the usage representation is useful, we trained each of our core models, plus DNN, on both the API usage and frequency datasets. The results are shown in Tables 11 and 9 respectively. Two things are notable: first, the differences in model accuracies are relatively small. Second, the ranking of the models changes depending on the representation, with DNNs doing best on the frequency dataset and RFs doing best on the usage dataset. Perhaps coincidentally, this justifies the model choices of Ma et al.
3.4. Reducing the Number of API Calls
Our results indicate that using API call features to build ML models leads to very good rates of malware detection. However, the large number of API calls available in the current Android SDK results in a high computational overhead when building models. In this section, we look at ways of reducing the size of the feature set.
Jung et al. (Jung et al., 2018) approached this by analysing the top 50 API calls used in benign applications then malware applications, resulting in two feature sets, one from each of the applications type, which they used to train RF models. The approach was evaluated using a dataset of 30,159 benign and 30,084 malicious applications. The results of our reimplementation are shown in Table 10. Of the two feature sets, the use of the top 50 malware API calls leads to better models, and this is supported by a Kruskal-Wallis test, which showed a significant difference (H = 14.23, p $<$ 0.05). However, the rates of accuracy achieved in our experiments are considerably less than those reported by Jung et al, and the resulting models performed relatively poorly. This difference may in part be due to the much larger size of the API call set in contemporary versions of Android.
Table 10. Result comparison between Jung et al. (Jung et al., 2018) and our reimplementation, showing the performance of models trained just using the top 50 API calls found in malware and the top 50 API calls found in benign software
Jung et al. Results Our Results
Dataset Accuracy Precision F1-Score TPR TNR Accuracy Precision F1-Score TPR TNR
Top 50 Benign 0.999 - - - - 0.701±$\pm 0.002$ 0.677±$\pm 0.003$ 0.721±$\pm 0.003$ 0.632±$\pm 0.005$ 0.771±$\pm 0.004$
Top 50 Malware 0.978 - - - - 0.762±$\pm 0.002$ 0.691±$\pm 0.002$ 0.799±$\pm 0.001$ 0.948±$\pm 0.002$ 0.576±$\pm 0.002$
Table 11. Result comparison between Muzaffar et al. (Muzaffar et al., 2021) and our reimplementation, showing the relative performance of different ML models trained using the full API usage feature set
Muzaffar et al. Results Our Results
Classifier Accuracy Precision F1-Score TPR TNR Accuracy Precision F1-Score TPR TNR
SVM 0.955±$\pm 0.002$ 0.957±$\pm 0.004$ 0.955±$\pm 0.002$ 0.953±$\pm 0.004$ 0.957±$\pm 0.004$ 0.953±$\pm 0.004$ 0.957±$\pm 0.002$ 0.954±$\pm 0.001$ 0.956±$\pm 0.003$ 0.958±$\pm 0.001$
Random Forest 0.959±$\pm 0.001$ 0.960±$\pm 0.001$ 0.959±$\pm 0.002$ 0.957±$\pm 0.001$ 0.960±$\pm 0.002$ 0.966±$\pm 0.002$ 0.968±$\pm 0.002$ 0.966±$\pm 0.002$ 0.964±$\pm 0.003$ 0.969±$\pm 0.001$
DT 0.940±$\pm 0.000$ 0.938±$\pm 0.003$ 0.941±$\pm 0.002$ 0.943±$\pm 0.002$ 0.938±$\pm 0.002$ 0.945±$\pm 0.002$ 0.950±$\pm 0.002$ 0.954±$\pm 0.001$ 0.947±$\pm 0.002$ 0.950±$\pm 0.002$
Naïve Bayes 0.744±$\pm 0.002$ 0.673±$\pm 0.003$ 0.789±$\pm 0.002$ 0.957±$\pm 0.002$ 0.531±$\pm 0.005$ 0.750±$\pm 0.001$ 0.675±$\pm 0.001$ 0.791±$\pm 0.002$ 0.957±$\pm 0.002$ 0.541±$\pm 0.003$
AdaBoost 0.943±$\pm 0.002$ 0.943±$\pm 0.001$ 0.943±$\pm 0.002$ 0.944±$\pm 0.004$ 0.942±$\pm 0.001$ 0.943±$\pm 0.002$ 0.943±$\pm 0.003$ 0.943±$\pm 0.002$ 0.944±$\pm 0.004$ 0.943±$\pm 0.001$
kNN - - - - - 0.942±$\pm 0.003$ 0.942±$\pm 0.002$ 0.944±$\pm 0.002$ 0.941±$\pm 0.002$ 0.944±$\pm 0.002$
DNN (16) - - - - - 0.961±$\pm 0.001$ 0.962±$\pm 0.003$ 0.963±$\pm 0.002$ 0.959±$\pm 0.001$ 0.962±$\pm 0.001$
Table 12. Result comparison between Muzaffar et al. (Muzaffar et al., 2021) and our reimplementation, showing the relative performance of random forest models trained using API usage feature sets reduced using different feature selection algorithms
Muzaffar et al. Results Our Results
Features Features Accuracy Precision F1-Score TPR TNR Features Accuracy Precision F1-Score TPR TNR
MI 10,000 0.962±$\pm 0.001$ 0.960±$\pm 0.001$ 0.961±$\pm 0.002$ 0.959±$\pm 0.002$ 0.963±$\pm 0.002$ 10,000 0.967±$\pm 0.001$ 0.969±$\pm 0.001$ 0.967±$\pm 0.002$ 0.963±$\pm 0.002$ 0.970±$\pm 0.002$
VT 6,443 0.961±$\pm 0.002$ 0.964±$\pm 0.001$ 0.961±$\pm 0.002$ 0.958±$\pm 0.002$ 0.964±$\pm 0.001$ 6,443 0.968±$\pm 0.002$ 0.971±$\pm 0.002$ 0.968±$\pm 0.002$ 0.965±$\pm 0.002$ 0.971±$\pm 0.001$
PCC 15,000 0.951±$\pm 0.001$ 0.952±$\pm 0.002$ 0.951±$\pm 0.002$ 0.938±$\pm 0.001$ 0.965±$\pm 0.002$ 15,000 0.961±$\pm 0.001$ 0.970±$\pm 0.003$ 0.961±$\pm 0.002$ 0.950±$\pm 0.001$ 0.971±$\pm 0.002$
Chi-Square - - - - - - 10,000 0.961±$\pm 0.003$ 0.961±$\pm 0.004$ 0.954±$\pm 0.002$ 0.967±$\pm 0.001$ 0.953±$\pm 0.001$
Figure 4. Relationship between number of API calls used and model accuracy for API usage features ranked using various methods
Figure 5. Number of API usage features selected according to variance threshold
Although these results suggest that 50 API calls are insufficient to train accurate models, from a practical perspective, it remains desirable to work with smaller feature sets. Consequently, in Muzaffar et al. (Muzaffar et al., 2021), we investigated the use of dimensionality reduction in API feature sets more broadly. We used a dataset of 20,000 benign applications and 20,000 malicious applications, and our results showed that the accuracy of the ML models could be improved by reducing the number of API calls by approximately 95% from the original set of 134,207. We trained SVM, RF, DT, AdaBoost and Naïve Bayes models, and used MI, VT and PCC for feature selection.
We reimplemented the study by using the current, considerably larger, dataset of 62,000 benign and 62,000 malicious applications. Table 11 shows the results for the full API calls dataset. We also added kNN and Chi-square and compared the results. The RF models achieved the best results, followed by SVM. Naïve Bayes was included because of its training efficiency with large feature sets, but it performed poorly in terms of discrimination.
The feature selection algorithms were then used to reduce the feature set sizes to between 100 and 30,000 API calls. Fig. 4 shows the results for model accuracy, indicating that both the feature set size and the feature selection algorithm used have an impact on performance. In each case, it can be seen that accuracy plateaus beyond a certain number of API calls, supporting the idea that we only need to use a subset of the full Android API when training malware detection models.
Regardless of feature set size, RFs were the best models. Table 12 shows the RF performance metrics for each of the feature selection algorithms. A Kruskal-Wallis test between the best performing RF model trained on the full API set and the models from Table 12 indicated significant differences (H = 32.59, p $<$ 0.05). Post hoc Dunn’s tests then showed that feature sets derived using variance threshold led to significantly better models. Notably, these models could outperform the full API set models across all metrics while using only 5% (6,443) of the total API calls.
Fig. 5 shows variance values for feature set sizes of 100 to 30,000. This shows that about half the features have a variance below 0.025, which also signifies that many of the API calls are unlikely to play a useful role in discrimination. Overall, these results show that the number of API calls can be significantly reduced without loss of accuracy.
3.5. The Drebin feature set
The Drebin dataset of Arp et al. (Arp et al., 2014) is widely used in Android malware detection, both due to its availability and the large number of extracted features. They provided the following features for each application in the dataset:
  • Hardware components: The set of hardware components that an application requests.
  • Requested permissions: The set of permissions an application requests in their manifest file.
  • App components: The different components the application requests in the manifest.
  • Filtered intents: Intents used by the application; these could include intents like “BOOT_COMPLETED”. Applications use intents for inter-process communication.
  • Restricted API calls: Some critical calls are restricted by Android, these are looked for in the Dex files.
  • Used permissions: Felt et al. (Felt et al., 2011) introduced a method which was used to match API calls to permissions, hence obtaining the permissions which are actually used by the application.
  • Suspicious API calls: API calls that can access sensitive data.
  • Network addresses: These include IP addresses, URLs, and hostnames.
To get an indication of whether this feature set remains useful within a current Android environment, we extracted the same features from our set of contemporary applications. Table 13 outlines the differences in the datasets, showing a growth from 535,000 to three million distinct features. Table 13 also shows the accuracy of SVM models trained on these two datasets. This shows that, although not as effective as the API call models we looked at earlier, the Drebin feature set does remain competitive.
Table 13. Comparison of samples and Drebin feature counts in Arp et al’s original dataset and our dataset, also showing the accuracy of SVM models trained on these two datasets
Benign Malware No. of features Accuracy
Drebin 123,453 5,560 535,000 0.94
Our Dataset 62,000 62,000 3 million reduced to 600,000 0.956
Table 14. Comparison of models trained on reduced Drebin feature sets, using different feature selection algorithms, showing results for the most effective ML model in each case
Feature Selection Classifier Number of Features Accuracy F1-Score Precision TPR TNR
Features used twice SVM 600,000 0.956±$\pm 0.002$ 0.952±$\pm 0.003$ 0.949±$\pm 0.002$ 0.961±$\pm 0.002$ 0.948±$\pm 0.001$
MI SVM 30000 0.897±$\pm 0.002$ 0.895±$\pm 0.003$ 0.887±$\pm 0.003$ 0.897±$\pm 0.003$ 0.891±$\pm 0.001$
VT SVM 35000 0.893±$\pm 0.002$ 0.893±$\pm 0.002$ 0.886±$\pm 0.004$ 0.894±$\pm 0.002$ 0.886±$\pm 0.002$
Chi-square RF 25000 0.895±$\pm 0.002$ 0.892±$\pm 0.004$ 0.889±$\pm 0.003$ 0.896±$\pm 0.002$ 0.889±$\pm 0.002$
PCC RF 35000 0.891±$\pm 0.002$ 0.890±$\pm 0.002$ 0.885±$\pm 0.002$ 0.894±$\pm 0.002$ 0.884±$\pm 0.003$
However, the size of the feature set means that it is computationally very challenging to train models. To address this, we only used the features that were used by at least two applications in the dataset, thereby reducing the feature set to 600,000. This allowed us to construct a dataset of a similar dimensionality to the original one. We then retrained the SVM models on this reduced dataset, and Table 14 shows that this led to a small reduction in accuracy.
We then looked at whether feature selection algorithms could be used to reduce the Drebin feature set to an even smaller set of useful features. Table 14 shows the results of applying the core set of feature selection algorithms and ML models. Although this significantly decreases the feature set size, it also leads to a significant decrease in accuracy. To conclude, the Drebin feature set does perform well; however, it may be difficult to use it in a production environment and reducing the number of features using feature selection algorithms decreases the performance of the models significantly.
3.6. Opcodes
Rather than a sequence of API calls, a program can instead be viewed, at a lower level, as a sequence of machine language instructions, or opcodes. A relatively small number of studies have considered opcode analysis as a basis for Android malware detection, and most have involved extracting features in the form of n-grams of opcodes, referred to as n-opcodes. For example, lets say that an application $A$ is the sequence $op_{1}$, $op_{2}$, $op_{3}$, $op_{4}$. A 2-gram, also known as bigram, representation of $A$ would be three bigrams, namely [$op_{1}$, $op_{2}$], [$op_{2}$, $op_{3}$] and [$op_{3}$, $op_{4}$]. To measure the benefits of this approach, and compare it against more traditional static features like permissions and API calls, we reimplement a study by Kang et al. (Kang et al., 2016), which considered both usage and frequency representations (see Section 3.3) of n-opcodes.
Kang et al. used a dataset comprising 1,260 benign and 1,260 malware applications. They extracted n-opcodes of sizes one up to 10, and then used MI to reduce the number of features, before training NB, RF, SVM and partial decision tree (PART) models. In our reimplementation, we replaced PART with our standard DT model and also trained kNNs. We used our four core feature selectors. Table 15 shows the number of n-opcodes for different values of $n$ up to 10 both for our dataset and the dataset used by Kang et al. This indicates a significant increase in the number of unique n-opcodes, particularly for larger values of $n$, making feature selection a necessary step. Table 16 shows the number of features for each value of $n$ after carrying out feature selection using MI, which we found to be the most effective feature selector in our experiments.
Table 15. Number of unique n-opcodes for different values of $n$
Kang et al. Our Dataset
n n-opcodes n n-opcodes
1 214 1 218
2 22,371 2 44,411
3 399,598 3 1,049,741
4 2,201,377 4 8,474,254
5 6,458,246 5 35,033,022
6 12,969,857 6 95,082,086
7 20,404,473 7 195,415,174
8 27,366,890 8 330,418,164
9 33,024,116 9 484,711,644
Table 16. Number of selected n-opcodes using mutual information
n Usage Frequency
1 99 159
2 3,991 5,387
3 14,321 20,219
4 38,456 45,201
5 46,781 49,246
6 53,578 56,983
7 60,210 64,392
8 58,219 62,239
9 57,985 59,238
10 58,456 59,249
Figure 6. Accuracy of ML models trained on usage and frequency-based $n$-opcode features selected using mutual information, showing the effect of changing the value of $n$
Fig. 6 shows the accuracy of the models trained on the features selected by MI, and also shows the effect of varying $n$. In general, there appears to be little benefit to increasing $n$ much beyond 4, especially given the lower computational effort associated with smaller n-opcodes. Among the four ML models, it can be seen that RF models performed best for all values of $n$ and both feature representations. The average accuracies for models that use usage features were slightly higher than those which used frequency features, but the difference was small in comparison to the effect of ML model choice.
Table 17 summarises the performance of the best models. A Kruskal-Wallis test applied to these results indicates a significant difference (H = 36.44, p $<$ 0.05) between models, with Dunn’s tests showing that RF and SVM models are significantly better than the others. The performance of our models was broadly similar to those reported by Kang et al. Notably, these models have similar performance to models based on API calls. Opcodes take less time to extract compared to API calls and are much easier to train computationally on the base feature sets. However, it is much easier to interpret the API call results as APIs correspond to specific features in the Android SDK.
Table 17. Result comparison between Kang et al. (Kang et al., 2016) and our reimplementation, showing the effect of ML model choice when using usage and frequency-based $n$-opcode features, and also showing the optimal value of $n$ in each case. F1-scores for the original study are approximate values as they were read from plots provided by the authors.
Kang et al. Our Results
Classifier n Accuracy Precision F1-Score TPR TNR Classifier n Accuracy Precision F1-Score TPR TNR
Usage Usage
SVM 3 - - 0.98 - - SVM 4 0.968±$\pm 0.002$ 0.971±$\pm 0.001$ 0.971±$\pm 0.002$ 0.961±$\pm 0.001$ 0.971±$\pm 0.002$
Random Forest 4 - - 0.98 - - Random Forest 5 0.973±$\pm 0.002$ 0.979±$\pm 0.002$ 0.975±$\pm 0.001$ 0.965±$\pm 0.001$ 0.979±$\pm 0.002$
Decision Trees 5 - - 0.98 - - Decision Trees 5 0.952±$\pm 0.001$ 0.952±$\pm 0.001$ 0.951±$\pm 0.003$ 0.949±$\pm 0.002$ 0.952±$\pm 0.001$
Naïve Bayes 3 - - 0.84 - - Naïve Bayes 5 0.832±$\pm 0.002$ 0.834±$\pm 0.002$ 0.828±$\pm 0.001$ 0.821±$\pm 0.003$ 0.838±$\pm 0.002$
Frequency Frequency
SVM 4 - - 0.96 - - SVM 5 0.956±$\pm 0.001$ 0.954±$\pm 0.002$ 0.956±$\pm 0.002$ 0.958±$\pm 0.003$ 0.954±$\pm 0.003$
Random Forest 8 - - 0.97 - - Random Forest 5 0.965±$\pm 0.002$ 0.969±$\pm 0.003$ 0.967±$\pm 0.002$ 0.961±$\pm 0.002$ 0.970±$\pm 0.002$
Decision Trees 3 - - 0.97 - - Decision Trees 6 0.940±$\pm 0.001$ 0.942±$\pm 0.001$ 0.937±$\pm 0.002$ 0.935±$\pm 0.003$ 0.941±$\pm 0.001$
Naïve Bayes 4 - - 0.85 - - Naïve Bayes 5 0.791±$\pm 0.002$ 0.794±$\pm 0.001$ 0.784±$\pm 0.003$ 0.781±$\pm 0.002$ 0.798±$\pm 0.001$
A common approach within deep learning is to convert numerical data into pseudoimages and use these to train a convolutional neural network (CNN). An example of this approach within the Android malware detection literature can be found in Xiao and Yang (Xiao and Yang, 2019), who converted disassembled opcode files into RGB images. To determine whether this has any advantage over n-opcode encodings, we reimplemented the same approach and wrote an algorithm that converts opcode sequence files into RGB images. We then used these to train a CNN model, using the same CNN topology as Xiao and Yang, which comprises standard convolutional, pooling and fully-connected layers. Xiao and Yang trained and evaluated their model using a dataset of 4,406 benign and 6,134 malicious applications. Table 18 compares their results against those of our reimplementation, which reported slightly lower metric scores. These shows that there is generally no benefit to using this more complex modelling approach over the simpler models reported in Table 17. This is supported by a statistical analysis, with Kruskal-Wallis (H = 56.44, p $<$ 0.05) and Dunn’s test showing Xiao and Yang’s approach to be significantly worse on our dataset than the best models in Table 17.
Table 18. Result comparison between Xiao and Yang’s (Xiao and Yang, 2019) CNN model trained on image-based opcode features and our reimplementation
Xiao and Yang’s Results Our Results
Accuracy Precision F1-Score TPR TNR Accuracy Precision F1-Score TPR TNR
0.93 0.936/0.921 0.94 0.944/0.910 0.903 0.913±$\pm 0.003$ 0.912±$\pm 0.001$ 0.916±$\pm 0.003$ 0.921±$\pm 0.001$ 0.911±$\pm 0.001$
Another example of using deep learning for opcode analysis can be found in a study by Yeboah and Baz Musah (Yeboah and Baz Musah, 2022). In this case, they used a 1D CNN; that is, a one-dimensional CNN which can be applied to sequential numeric data. A notable element of their approach was the use of a word2vec embedding model to represent each opcode as a single numeric value, allowing opcode sequences to be represented as dense real-valued vectors (as opposed to using one-hot encodings). The embedding model was trained separately on unigram and bigram opcode sequences, resulting in matrices of size M×N, where M is the number of unique opcodes and N is the vector dimension. N was set to 64 for both unigram and bigram embeddings. The 1D CNN comprised convolutional layers, pooling layers, and fully connected layers. Table 19 shows the results of our reimplementation, and again shows that there is no benefit to using a more complex modelling approach over the simpler models reported in Table 17. Our results were significantly lower than those reported by Yeboah and Baz Musah, who used a small (4948 malware, 2477 benign) though relatively recent dataset.
Table 19. Result comparison between Yeboah and Baz Musah’s (Yeboah and Baz Musah, 2022) 1D CNN model trained on sequence-based opcode features and our reimplementation
Yeboah and Baz Musah’s Results Our Results
Accuracy Precision F1-Score TPR TNR Accuracy Precision F1-Score TPR TNR
- 0.98 0.97 0.97 - 0.936±$\pm 0.001$ 0.933±$\pm 0.002$ 0.935±$\pm 0.002$ 0.923±$\pm 0.002$ 0.929±$\pm 0.001$
3.7. Comparison of static modelling approaches
To identify the best combination of features, ML models and feature selection algorithms for static analysis, we carried out a Kruskal-Wallis test followed by a pairwise comparison between the modelling approaches from this section. The results indicate that five of the models have statistically better accuracies than the others. Three of these are API call models: SVMs trained using 10,000 features selected using Chi-square, RF models trained using 10,000 features selected using MI, and RF models trained using 6,443 features selected using variance threshold (Table 12). The others are the RF usage 5-opcode model, trained using 46,781 features selected using mutual information (Table 17) and DNN trained using API frequency (Table 7). Neither of these five modelling approaches appears to be significantly better than the others. This suggests that the best classes of features to use for static analysis are opcodes and API calls, and the best ML model to use is likely to be RF. However, the best choice of feature selection algorithm depends on the feature type and model type.
4. Dynamic Analysis
In this section, we focus on the use of dynamic features in building Android malware detection models. We reimplement five previous studies, and also report the results of new experiments that are intended to fill knowledge gaps found in these studies.
Dynamic analysis is not as common as static analysis. This is due to several factors, including the complex set up, relatively high computational costs, and the longer time required to run the analysis. A particular issue when running dynamic analysis is that each application needs to be executed on an emulator in order to record its dynamic features. During this process, some applications terminate unexpectedly. In some cases, the emulator itself crashes. Taking this into account, the set of applications for which we were able to complete dynamic analysis comprises 53,960 benign and 53,202 malware. However, this remains the largest and most up-to-date dataset used for dynamic analysis, to the best of our knowledge.
4.1. System Calls
System calls are the most commonly used dynamic features in the literature. Android applications use system calls to communicate with the kernel of the OS, and the strace module can be used to trace system calls that are made by an application during its runtime.
Ananya et al. (Ananya et al., 2020) used sequences of system calls represented as unigrams, bigrams, and trigrams as the features of their models. They also used feature selection to select relevant features, and trained four ML models: linear regression, decision tree, RF and XGBoost. They used a dataset of 2,475 benign applications and 2,474 malware. The authors proposed two novel approaches to feature selection. The first, called SAILS, builds on the conventional feature selection algorithms mutual information, Chi-square and DFS. These are used to score the benign and malware features. SAILS then sorts the resulting scores in descending order, to create two sorted lists of benign and malware features. It then creates a final list that is the union of these two lists. The second feature selection algorithm proposed was WFS, which assigns weights to system calls. This is done by calculating the ratio of the number of occurrences of the system call in malware to the number of occurrences in the sum of benign and malware. In all the experiments, the authors used SAILS, WFS, mutual information, Chi-square and DFS to select relevant features. In our reimplementation, we also add PCC.
Table 20. Result comparison between Ananya et al. (Ananya et al., 2020) and our reimplementation, showing the effect of ML model choice when trained using system call features represented as unigrams, bigrams and trigrams. The most effective feature selection method is shown in each case.
Ananya et al.  Results Our Results
Unigram
Classifier Feature Selection Accuracy Precision F1-Score TPR TNR Feature Selection Accuracy Precision F1-Score TPR TNR
SVM - - - - - - Mutual information 0.928±$\pm 0.001$ 0.926±$\pm 0.001$ 0.924±$\pm 0.003$ 0.922±$\pm 0.002$ 0.929±$\pm 0.001$
Linear Regression Mutual information 0.977 - - - - Mutual information 0.921±$\pm 0.002$ 0.922±$\pm 0.001$ 0.919±$\pm 0.002$ 0.915±$\pm 0.001$ 0.923±$\pm 0.002$
Random Forest WFS 0.972 - - - - Mutual information 0.934±$\pm 0.001$ 0.935±$\pm 0.001$ 0.933±$\pm 0.003$ 0.931±$\pm 0.003$ 0.936±$\pm 0.001$
Decision Tree Mutual information 0.965 - - - - Mutual information 0.925±$\pm 0.002$ 0.926±$\pm 0.002$ 0.921±$\pm 0.002$ 0.922±$\pm 0.001$ 0.926±$\pm 0.001$
XGBoost WFS 0.966 - - - - Mutual information 0.933±$\pm 0.001$ 0.931±$\pm 0.001$ 0.930±$\pm 0.002$ 0.931±$\pm 0.002$ 0.932±$\pm 0.001$
kNN - - - - - - Mutual information 0.922±$\pm 0.002$ 0.918±$\pm 0.002$ 0.921±$\pm 0.002$ 0.917±$\pm 0.003$ 0.923±$\pm 0.002$
Bigram
Classifier Feature Selection Accuracy Precision F1-Score TPR TNR Feature Selection Accuracy Precision F1-Score TPR TNR
SVM - - - - - - Mutual information 0.941±$\pm 0.002$ 0.939±$\pm 0.002$ 0.942±$\pm 0.003$ 0.938±$\pm 0.002$ 0.943±$\pm 0.001$
Linear Regression Chi-square 0.991 - - - - Mutual information 0.936±$\pm 0.001$ 0.930±$\pm 0.001$ 0.934±$\pm 0.002$ 0.938±$\pm 0.002$ 0.931±$\pm 0.001$
Random Forest Chi-square 0.988 - - - - Chi-square 0.945±$\pm 0.002$ 0.941±$\pm 0.001$ 0.943±$\pm 0.001$ 0.945±$\pm 0.002$ 0.941±$\pm 0.002$
Decision Tree Chi-square 0.977 - - - - Mutual information 0.931±$\pm 0.003$ 0.929±$\pm 0.001$ 0.929±$\pm 0.001$ 0.931±$\pm 0.002$ 0.929±$\pm 0.003$
XGBoost DFS 0.994 - - - - Chi-square 0.943±$\pm 0.002$ 0.943±$\pm 0.001$ 0.942±$\pm 0.003$ 0.939±$\pm 0.002$ 0.944±$\pm 0.002$
kNN - - - - - - Chi-square 0.932±$\pm 0.001$ 0.929±$\pm 0.003$ 0.931±$\pm 0.001$ 0.926±$\pm 0.002$ 0.931±$\pm 0.002$
Trigram
Classifier Feature Selection Accuracy Precision F1-Score TPR TNR Feature Selection Accuracy Precision F1-Score TPR TNR
SVM - - - - - - Chi-square 0.947±$\pm 0.001$ 0.945±$\pm 0.002$ 0.946±$\pm 0.002$ 0.947±$\pm 0.002$ 0.951±$\pm 0.003$
Linear Regression Mutual information 0.994 - - - - Mutual information 0.932±$\pm 0.003$ 0.932±$\pm 0.001$ 0.934±$\pm 0.003$ 0.937±$\pm 0.002$ 0.932±$\pm 0.002$
Random Forest Mutual information 0.982 - - - - Chi-square 0.951±$\pm 0.002$ 0.951±$\pm 0.001$ 0.950±$\pm 0.002$ 0.949±$\pm 0.003$ 0.952±$\pm 0.001$
Decision Tree WFS 0.987 - - - - Mutual information 0.945±$\pm 0.003$ 0.943±$\pm 0.001$ 0.942±$\pm 0.002$ 0.943±$\pm 0.002$ 0.943±$\pm 0.001$
XGBoost Mutual information 0.992 - - - - Chi-square 0.950±$\pm 0.002$ 0.948±$\pm 0.001$ 0.947±$\pm 0.002$ 0.947±$\pm 0.002$ 0.948±$\pm 0.002$
kNN - - - - - - Chi-square 0.942±$\pm 0.002$ 0.941±$\pm 0.001$ 0.943±$\pm 0.002$ 0.938±$\pm 0.001$ 0.941±$\pm 0.002$
Table 20 shows the best results for each ML model using unigrams, bigrams, and trigrams, and also indicates the most effective feature selection approach in each case. Notably, the performance metrics achieved in our reimplementation are substantially lower than those published in the original study. Ananya et al. reported accuracies of above 99%, whereas our best models achieved an accuracy of around 95%. This shows the importance of reimplementing studies using larger contemporary datasets, since it leads to the conclusion that dynamic analysis using system calls is less effective than the less expensive static analysis approaches considered earlier.
Neverthelesss, our results suggest that larger n-grams are more effective, and that RFs are again the best performing classifiers. A Kruskal-Wallis test between the models in Table 20 showed that there is a statistically significant difference in accuracy (H = 46.64, p $<$ 0.05). Post hoc Dunn’s tests confirmed that RFs outperformed the decision trees and linear regression models. Mutual information and Chi-square selected the best features for malware detection. Fig. 7 shows the accuracy rates of the ML models with different number of features selected by mutual information and Chi-square. The accuracy does not increase significantly, or in some cases dropped, after 90 features for unigrams, 5,000 features for bigrams using mutual information, 7,000 features for bigrams using Chi-square, and 80,000 features for trigrams.
(a) unigram
(b) bigram
(c) trigram
Figure 7. Mean accuracy of ML models trained on system call features represented as unigrams, bigrams and trigrams when using mutual information and Chi-square to select features, showing the effect of varying the feature count n.
Table 21. Result comparison between Malik et al. (Malik et al., 2019) and our reimplementation, showing the performance of kNN models trained on system call frequency features and LSTM models trained on system call sequences
Malik et al. Results Our Results
Classifier Accuracy Precision F1-Score TPR TNR Classifier Accuracy Precision F1-Score TPR TNR
kNN (3) - 0.852 0.846 0.839 - kNN (10) 0.820±$\pm 0.003$ 0.869±$\pm 0.002$ 0.806±$\pm 0.002$ 0.752±$\pm 0.002$ 0.887±$\pm 0.004$
LSTM (128) - 0.786 0.859 0.946 - LSTM (128) 0.7521±$\pm 0.001$ 0.753±$\pm 0.003$ 0.749±$\pm 0.002$ 0.732±$\pm 0.002$ 0.758 ±$\pm 0.001$
Table 22. Comparison of core ML models trained on usage and frequency-based system call features
Usage Frequency
Classifier Accuracy Precision F1-Score TPR TNR Accuracy Precision F1-Score TPR TNR
SVM 0.830±$\pm 0.002$ 0.843±$\pm 0.002$ 0.827±$\pm 0.002$ 0.811±$\pm 0.003$ 0.849±$\pm 0.003$ 0.821±$\pm 0.002$ 0.823±$\pm 0.004$ 0.824±$\pm 0.002$ 0.810±$\pm 0.003$ 0.832±$\pm 0.003$
Random Forest 0.845±$\pm 0.002$ 0.858±$\pm 0.001$ 0.842±$\pm 0.002$ 0.826±$\pm 0.002$ 0.863±$\pm 0.001$ 0.839±$\pm 0.002$ 0.835±$\pm 0.002$ 0.833±$\pm 0.002$ 0.856±$\pm 0.002$ 0.821±$\pm 0.002$
DT 0.812±$\pm 0.003$ 0.818±$\pm 0.002$ 0.810±$\pm 0.002$ 0.803±$\pm 0.002$ 0.821±$\pm 0.001$ 0.802±$\pm 0.002$ 0.809±$\pm 0.003$ 0.806±$\pm 0.003$ 0.802±$\pm 0.001$ 0.815±$\pm 0.003$
kNN 0.820±$\pm 0.003$ 0.869±$\pm 0.002$ 0.806±$\pm 0.002$ 0.752±$\pm 0.002$ 0.887±$\pm 0.004$ 0.823±$\pm 0.002$ 0.826±$\pm 0.003$ 0.828±$\pm 0.003$ 0.822±$\pm 0.001$ 0.828±$\pm 0.001$
We reimplemented one more study that built models solely from system call features, carried out by Malik et al. (Malik et al., 2019). This study is notable for using a different group of ML models and a different representation of the features. In particular, they used a 3-layer LSTM to classify malware based on the complete sequence of system calls used by an application, comparing this against a kNN model trained on a system call usage feature set. Both $k$ and the number of neurons per LSTM layer were optimised. Table 21 shows the results. Like Malik et al., we found that there was no benefit to using a more complex, slower to train, sequence-based model, since the best kNN model (with $k=3$) substantially out-performed the best LSTM (with 128 neurons per layer). However, our LSTM model performed much better than the model reported by Malik et al.
To get a slightly broader perspective, we also trained our standard set of ML models using both the system call usage set and the corresponding system call frequency feature set. The results are shown in Table 22. Reflecting our previous findings, the effect of moving between a usage-based and a frequency-based representation is relatively small — although, generally, the best models were found when using the usage-based feature set. Increasing the model set shows that kNN is not the best choice, with RF models leading to the best accuracy and F1-scores for both feature representations. This was confirmed by Kruskal-Wallis (H = 22.41, p $<$ 0.05) and Dunn’s tests, which showed RF to be statistically better than the other models, with no significant difference between RF usage and frequency models.
4.2. API Calls
Perhaps an obvious question at this point is whether API calls, which we found to be the most effective class of features in models resulting from static analysis, would also be beneficial for dynamic analysis. Afonso et al. (Afonso et al., 2015) addressed this question to a certain extent by using information about both system calls and API calls in their models. Specifically, they used call frequencies as features, and a dataset of 2,968 benign and 4,552 malicious applications. We reimplemented their approach using SVM, RF, decision trees, kNN and Naïve Bayes. Afonso et al. also considered several other models, but those performed poorly in their study, so we did not reimplement them.
Table 23. Result comparison between Afonso et al. (Afonso et al., 2015) and our reimplementation, showing performance of ML models trained on a combined system and API call feature set
Afonso et al. Results Our Results
Classifier Accuracy Precision F1-Score TPR TNR Classifier Accuracy Precision F1-Score TPR TNR
Random Forest 0.968 0.975 0.968 0.961 0.976 Random Forest 0.962±$\pm 0.002$ 0.968±$\pm 0.002$ 0.964±$\pm 0.001$ 0.959±$\pm 0.003$ 0.969±$\pm 0.002$
SVM 0.951±$\pm 0.001$ 0.953±$\pm 0.001$ 0.952±$\pm 0.001$ 0.949±$\pm 0.002$ 0.959±$\pm 0.002$
kNN 0.946±$\pm 0.003$ 0.943±$\pm 0.003$ 0.946±$\pm 0.003$ 0.943±$\pm 0.003$ 0.948±$\pm 0.002$
DT 0.936±$\pm 0.001$ 0.938±$\pm 0.001$ 0.935±$\pm 0.001$ 0.933±$\pm 0.003$ 0.938±$\pm 0.002$
Naïve Bayes 0.857±$\pm 0.002$ 0.860±$\pm 0.002$ 0.856±$\pm 0.003$ 0.851±$\pm 0.002$ 0.862±$\pm 0.002$
Table 24. Performance of RF models trained only on API calls, also comparing the benefit of usage and frequency-based features
Usage Frequency
Features Accuracy Precision F1-Score TPR TNR Accuracy Precision F1-Score TPR TNR
Full Set 0.829±$\pm 0.002$ 0.839±$\pm 0.003$ 0.826±$\pm 0.003$ 0.813±$\pm 0.003$ 0.844±$\pm 0.001$ 0.812±$\pm 0.002$ 0.815±$\pm 0.001$ 0.812±$\pm 0.002$ 0.802±$\pm 0.002$ 0.831±$\pm 0.001$
MI 0.855±$\pm 0.003$ 0.858±$\pm 0.002$ 0.851±$\pm 0.003$ 0.855±$\pm 0.001$ 0.845±$\pm 0.001$ 0.842±$\pm 0.003$ 0.841±$\pm 0.001$ 0.833±$\pm 0.003$ 0.816±$\pm 0.003$ 0.846±$\pm 0.003$
PCC 0.842±$\pm 0.002$ 0.843±$\pm 0.002$ 0.848±$\pm 0.002$ 0.834±$\pm 0.002$ 0.844±$\pm 0.003$ 0.846±$\pm 0.002$ 0.851±$\pm 0.002$ 0.855±$\pm 0.003$ 0.842±$\pm 0.002$ 0.833±$\pm 0.001$
VT 0.833±$\pm 0.002$ 0.845±$\pm 0.002$ 0.840±$\pm 0.002$ 0.845±$\pm 0.003$ 0.831±$\pm 0.002$ 0.845±$\pm 0.001$ 0.843±$\pm 0.001$ 0.843±$\pm 0.003$ 0.831±$\pm 0.003$ 0.829±$\pm 0.001$
Chi-square 0.829±$\pm 0.002$ 0.829±$\pm 0.002$ 0.806±$\pm 0.002$ 0.814±$\pm 0.002$ 0.834±$\pm 0.001$ 0.826±$\pm 0.003$ 0.821±$\pm 0.002$ 0.822±$\pm 0.002$ 0.804±$\pm 0.002$ 0.834±$\pm 0.002$
Table 23 shows the results of our reimplementation. The RF model does particularly well, and has comparable performance to the best static approaches. Notably, its performance is significantly higher than the system call-only models reported in the previous section (Kruksal-Wallis, H = 24.57, p $<$ 0.05, supported by Dunn’s tests). However, it should be noted those models used n-gram and sequence-based representations, rather than API call frequencies.
This still leaves the question of whether dynamic API calls alone can be used to build effective models. To address this, Table 24 shows the performance of RF models trained solely on API call frequencies. The results are shown for both the full feature set and for feature sets reduced using the core feature selection algorithms. It can be seen that the performance of these models is a lot lower than the models that used both system and API calls. This seems to suggest that whilst the combination of dynamic system calls and dynamic API calls is more effective than either of these alone, information about dynamic API calls is not sufficient to train good models. This is an interesting finding, given that information about static API calls was sufficient to train good models. The poor performance of using the dynamic API calls alone could also be explained by the fact that while all system calls are monitored, only a subset of the API calls can be monitored due to their large number (134,207).
We also show, in Table 24, the performance of models based upon usage (rather than frequency) of dynamic system and API calls. In common with our results from earlier studies, this shows that the choice of usage-based or frequency-based features has only a minor impact upon model performance. However, it is perhaps more notable within a dynamic context, where frequency-based features might be expected to contain information that were not available through static analysis alone, and highlights the fact that this additional information is not necessarily required in order to train good malware detection models.
4.3. Network Traffic
Table 25. Result comparison between Zulkifli et al. (Zulkifli et al., 2018) and our reimplementation, showing the performance of ML models trained using their TCP-based network traffic feature set
Zulkifli et al. Results Our Results
Dataset Classifier Accuracy F1-Score TPR TNR Classifier Accuracy Precision F1-Score TPR TNR
drebin DT 0.984 - 0.920 - SVM 0.961±$\pm 0.001$ 0.958±$\pm 0.002$ 0.957±$\pm 0.002$ 0.952±$\pm 0.002$ 0.961±$\pm 0.002$
contagiodumpset DT 0.976 - 0.920 - Random Forest 0.977±$\pm 0.003$ 0.969±$\pm 0.003$ 0.981±$\pm 0.003$ 0.978±$\pm 0.002$ 0.971±$\pm 0.002$
DT 0.973±$\pm 0.002$ 0.967±$\pm 0.003$ 0.977±$\pm 0.001$ 0.973±$\pm 0.002$ 0.969±$\pm 0.001$
kNN 0.961±$\pm 0.001$ 0.962±$\pm 0.003$ 0.959±$\pm 0.002$ 0.968±$\pm 0.001$ 0.962±$\pm 0.001$
Table 26. Result comparison between Wang et al. (Wang et al., 2016) and our reimplementation, showing the performance of ML models trained using their TCP and HTTP-based network traffic feature sets. For our reimplementation, we only show results of the best ML model for each feature set.
Wang et al. Results Our Results
Dataset Classifier Accuracy Precision F1-Score TPR TNR Classifier Accuracy Precision F1-Score TPR TNR
TCP DT 0.982 - - - - Random Forest 0.973±$\pm 0.002$ 0.967±$\pm 0.003$ 0.977±$\pm 0.001$ 0.973±$\pm 0.002$ 0.969±$\pm 0.001$
HTTP DT 0.997 - - - - Random Forest 0.963±$\pm 0.002$ 0.961±$\pm 0.002$ 0.965±$\pm 0.002$ 0.965±$\pm 0.002$ 0.961±$\pm 0.003$
Table 27. Performance of all core ML models trained on HTTP features or a combination of both TCP and HTTP features. Note that models trained only on TCP features are shown in Table 25.
Network Feature Classifier Accuracy Precision F1-Score TPR TNR
HTTP SVM 0.941±$\pm 0.003$ 0.942±$\pm 0.001$ 0.941±$\pm 0.002$ 0.943±$\pm 0.002$ 0.946±$\pm 0.003$
HTTP Random Forest 0.963±$\pm 0.002$ 0.961±$\pm 0.002$ 0.965±$\pm 0.002$ 0.965±$\pm 0.002$ 0.961±$\pm 0.003$
HTTP DT 0.932±$\pm 0.002$ 0.935±$\pm 0.003$ 0.937±$\pm 0.004$ 0.933±$\pm 0.001$ 0.932±$\pm 0.002$
HTTP kNN 0.951±$\pm 0.002$ 0.956±$\pm 0.003$ 0.955±$\pm 0.004$ 0.949±$\pm 0.003$ 0.959±$\pm 0.002$
TCP + HTTP SVM 0.938±$\pm 0.002$ 0.933±$\pm 0.003$ 0.934±$\pm 0.004$ 0.933±$\pm 0.001$ 0.932±$\pm 0.003$
TCP + HTTP Random Forest 0.952±$\pm 0.002$ 0.953±$\pm 0.002$ 0.956±$\pm 0.002$ 0.955±$\pm 0.004$ 0.951±$\pm 0.002$
TCP + HTTP DT 0.929±$\pm 0.002$ 0.926±$\pm 0.002$ 0.925±$\pm 0.003$ 0.922±$\pm 0.002$ 0.925±$\pm 0.003$
TCP + HTTP kNN 0.967±$\pm 0.004$ 0.964±$\pm 0.002$ 0.967±$\pm 0.001$ 0.969±$\pm 0.002$ 0.964±$\pm 0.003$
Dynamic analysis tools can also monitor network traffic during the execution of an application, and this provides another potentially useful source of information about the application’s behaviour. Previous works have focused on using network traffic features extracted from TCP and HTTP protocols.
Zulkifli et al. (Zulkifli et al., 2018) used DT with TCP-based features in their study. They used a small dataset of 500 benign and 200 malicious applications, the latter assembled from two existing malware collections (Drebin and contagiodumpset). Notably, all the applications were executed on a real Android device. They extracted the following features:
  • The average packet size during the runtime of the dynamic analysis
  • The average of the number of packets sent per flow between host and client during the runtime of the dynamic analysis
  • The average of the number of packets sent per flow between client and host during the runtime of the dynamic analysis
  • The average packet size in terms of bytes sent per flow during the runtime of the dynamic analysis
  • The average packet size in terms of bytes received per flow during the runtime of the dynamic analysis
  • Ratio of the number of incoming bytes to outgoing bytes of every packet during the runtime of the dynamic analysis
  • The average number of packets received per second during the course of the dynamic analysis
We extracted the same seven features for each of the applications in our dataset and trained the four core ML models. Table 25 shows the results, comparing our reimplementation with Zulkifli et al., who reported results separately for each of their two malware datasets. The best performing model was RF. Significantly, its accuracy and F1-score is the highest we have seen so far, suggesting that network traffic is an important source of information within a malware detection context.
Zulkifli et al. used a small dataset, presumably because they used a real Android device to perform their dynamic analysis. In this respect, it is interesting to see that our results show similar accuracy to the original paper. This supports the case for using an emulator to perform dynamic analysis, especially since an emulator is both a lot faster to configure and allows the OS image to be reloaded each time a dynamic analysis of an application is carried out, meaning that the analysis can be carried out in a controlled environment.
To further investigate network traffic features, we reimplemented another study by Wang et al. (Wang et al., 2016) involving network data. Wang et al. extracted both TCP and HTTP features from network traffic. Wang et al. used the same set of TCP features as Zulkifli et al., with the exception of average packet size. The features they extracted from HTTP headers were:
  • Host: This field specifies the host and port number that the request is sent to.
  • Request-URI: This is the uniform resource identifier from the requested source.
  • Request-Method: The action to be performed for a given resource. This includes GET and POST requests.
  • User-Agent: This holds information about the application, OS, vendor and version of the requesting user agent.
Wang et al. used a malware dataset consisting of 5,560 applications from the Drebin dataset and 8,312 benign applications downloaded from several application markets. The authors separately trained decision trees on both the HTTP and TCP datasets. In our reimplementation, we expanded on the original study by training our core set of ML models on the HTTP and TCP features both individually and in combination.
Table 26 shows the results, and compares them against those from the original study. For our reimplementation, Table 26 shows the results of the best ML models, which were RFs for both the HTTP and TCP feature sets. We observed no benefit to combining the two feature sets, but, for completeness, we also show results for the other models for both the HTTP and combined feature sets in Table 27. Whilst the results show good discrimination between benign applications and malware, we did not see the exceptionally high level of accuracy reported by Wang et al. for TCP features, and in fact found these to be less discriminative than HTTP features. These conclusions are supported by a positive Kruskal-Wallis test (H = 14.29, p $<$ 0.05) and Dunn’s tests.
Although network traffic can be used to train accurate models, it is worth noting that consistently generating network traffic in an application is a major challenge. For instance, a network connection issue may result in the application not generating any network traffic. Furthermore, depending on the coverage of the dynamic analysis, i.e. the different activities invoked during the analysis, this may impact the network traffic captured.
4.4. Comparison of dynamic modelling approaches
To identify the best combination of ML models and feature selection algorithm on dynamic features, we carried out a Kruskal-Wallis test followed by a pairwise comparison between the models from section 4. This showed that the network traffic models from Table 27 were significantly better than the others. However, as it can be a challenge to generate network traffic in dynamic analysis, it is useful to identify other well-performing dynamic approaches. Statistically, the best of the rest were the RF and XGBoost models trained on system call trigrams, selected using mutual information and Chi-square, respectively.
RF, as was the case for static features, was the best performing model for most of the dynamic features. Mutual information and Chi-square selected the most relevant features when used with both system calls and API calls. However, the results indicate that using static features generally produced the best results, with the notable exception of network traffic. System calls did lead to results comparable to the best performing static feature, API calls; however, the computational effort and time required to carry out a system call analysis probably outweighs the benefits.
5. Hybrid Analysis
Table 28. Result comparison between Kandukuru and Sharma (Kandukuru and Sharma, 2017) and our reimplementation, showing the performance of ML models trained on dynamic network traffic features and static permissions
Kandukuru and Sharma Results Our Results
Classifier Accuracy Precision F1-Score TPR TNR Classifier Accuracy Precision F1-Score TPR TNR
DT 0.956 - - 0.95 0.96 Random Forest 0.943±$\pm 0.001$ 0.941±$\pm 0.002$ 0.945±$\pm 0.002$ 0.948±$\pm 0.002$ 0.940±$\pm 0.002$
SVM 0.933±$\pm 0.002$ 0.936±$\pm 0.001$ 0.935±$\pm 0.002$ 0.939±$\pm 0.001$ 0.937±$\pm 0.003$
DT 0.921±$\pm 0.003$ 0.923±$\pm 0.003$ 0.923±$\pm 0.002$ 0.926±$\pm 0.003$ 0.925±$\pm 0.001$
kNN 0.936±$\pm 0.002$ 0.938±$\pm 0.002$ 0.939±$\pm 0.002$ 0.941±$\pm 0.004$ 0.937±$\pm 0.002$
Table 29. Reimplementation of Kapratwar et al.’s (Kapratwar et al., 2017) approach, in which models are trained on static permissions and dynamic system call features
Classifier Accuracy Precision F1-Score TPR TNR
SVM 0.937±$\pm 0.01$ 0.937±$\pm 0.02$ 0.937±$\pm 0.02$ 0.923±$\pm 0.01$ 0.951±$\pm 0.02$
Random Forest 0.948±$\pm 0.02$ 0.949±$\pm 0.01$ 0.946±$\pm 0.01$ 0.935±$\pm 0.01$ 0.959±$\pm 0.02$
Naïve Bayes 0.860±$\pm 0.02$ 0.861±$\pm 0.02$ 0.860±$\pm 0.02$ 0.831±$\pm 0.01$ 0.889±$\pm 0.01$
Logistic Regression 0.937±$\pm 0.02$ 0.938±$\pm 0.01$ 0.936±$\pm 0.02$ 0.925±$\pm 0.01$ 0.949±$\pm 0.02$
kNN 0.921±$\pm 0.02$ 0.924±$\pm 0.02$ 0.922±$\pm 0.03$ 0.886±$\pm 0.01$ 0.958±$\pm 0.01$
Decision Trees 0.919±$\pm 0.02$ 0.919±$\pm 0.02$ 0.917±$\pm 0.01$ 0.923±$\pm 0.02$ 0.917±$\pm 0.02$
Table 30. Reimplementation of Kapratwar et al.’s (Kapratwar et al., 2017) approach using reduced feature sets selected using mutual information
Classifier No. of Features Accuracy Precision F1-Score TPR TNR
SVM 300 0.936±$\pm 0.02$ 0.936±$\pm 0.01$ 0.935±$\pm 0.02$ 0.921±$\pm 0.01$ 0.951±$\pm 0.02$
Random Forest 250 0.948±$\pm 0.02$ 0.948±$\pm 0.03$ 0.949±$\pm 0.03$ 0.935±$\pm 0.01$ 0.961±$\pm 0.02$
Naive Bayes 100 0.865±$\pm 0.03$ 0.868±$\pm 0.02$ 0.864±$\pm 0.01$ 0.825±$\pm 0.01$ 0.907±$\pm 0.02$
Logistic Regression 300 0.936±$\pm 0.02$ 0.937±$\pm 0.02$ 0.935±$\pm 0.02$ 0.923±$\pm 0.02$ 0.949±$\pm 0.02$
kNN 150 0.922±$\pm 0.02$ 0.925±$\pm 0.01$ 0.922±$\pm 0.01$ 0.883±$\pm 0.03$ 0.962±$\pm 0.01$
DT 200 0.918±$\pm 0.02$ 0.917±$\pm 0.02$ 0.916±$\pm 0.02$ 0.920±$\pm 0.03$ 0.915±$\pm 0.01$
Hybrid analysis in theory uses the best of both worlds, static and dynamic features. As with dynamic analysis, hybrid studies are not as common as static analysis. In this section, we report and compare models that use hybrid features by reimplementing and extending past studies. As hybrid analysis uses dynamic features, in this section we use the slightly reduced dataset introduced in section 4, which consists of 53,960 benign and 53,202 malware.
We begin by reimplementing a study by Kandukuru and Sharma (Kandukuru and Sharma, 2017) that combined dynamic network traffic features with static permission features. The network traffic features included all seven of the TCP features listed in Section 4.3 and the authors used a malware dataset of 1,200 applications collected from 2010 to 2011 to train a decision tree model. In our reimplementation, we trained all four core ML models. Table 28 compares their performance with Kandukuru and Sharma’s results. Since their performance is worse than models trained only on network features (see Tables tables 25 and 26), this suggests that there is no benefit to including information about static permissions.
To further investigate the use of hybrid features, we reimplemented a study based around two frequently used groups of static and dynamic features, permissions and system calls. Kapratwar et al. (Kapratwar et al., 2017) carried out experiments using these features both separately and in combination. In the case of system calls, a frequency-based representation was used. For each experiment, the authors trained RF, decision trees, Naïve Bayes, logistic regression, kNN and SVM. However, as with most past hybrid works, Kapratwar et al. used a small dataset of just 103 malware and 97 benign applications. Therefore, a reimplementation of this approach using a larger contemporary dataset is essential to understand the effect of combining two of the most prevalent features in Android malware detection.
The results of our reimplementation are shown in Table 29, though note that we do not report the results of the original study in this case, since only AUC scores were provided. The performance of our models is comparable to the performance of models discussed earlier that used system calls or permissions separately, so again there does not appear to be any benefit to using a hybrid approach. We also used feature selection to select the most relevant features; mutual information was most useful for this, and Table 30 summarises the performance of the best models trained on the feature sets reduced according to mutual information. This suggests that there is a only a small benefit to doing this. However, as Fig. 8 shows, the number of features can be reduced by about a third without degrading performance. A Kruskal-Wallis test between the models in Table 30 showed that there was a statistically significant difference in accuracy (H = 25.62, p $<$ 0.05), and post hoc Dunn’s tests showed that the best performing ML model was RF.
Figure 8. Effect on ML model accuracies of reducing the number of system calls and permissions features using mutual information
6. Best-Performing Models
Table 31 lists the best classification models seen during this study. Two modelling approaches stand out in terms of the metric scores achieved when discriminating malware from benign applications. TCP network traffic models using random forests ($C_{9}$) achieved the highest scores in terms of accuracy, precision and TNR. Five-gram opcode models, again using random forests ($C_{6}$), achieved the highest scores in terms of F1 and TPR. As noted in Section 2.3, TPR is particularly important for malware detection, since a benign application labelled as malware is less likely to cause as many security problems as malware labelled as benign. It is also notable that 11 of the 12 classifiers listed in Table 31 use random forests, suggesting that this is a good default choice within malware detection. This reflects the wider success of tree-based ensemble models, especially when working with pre-extracted features (Grinsztajn et al., 2022).
There is also one kNN instance in the table ($C_{10}$), suggesting that there may sometimes be a benefit to using other models. However, we also observed that kNN is challenging in terms of scalability, since its inference time grows in line with the product of the number of samples and the number of features in the training data. This is an issue when using large feature sets such as API calls.
A notable omission within this table is the presence of any deep learning models. In the majority of Android malware detection studies, features are pre-extracted. In this situation, there is often little benefit to using a deep learning model (Grinsztajn et al., 2022), since their strength lies in their ability to extract features from raw data. Deep learning models also bring risks of overfitting, responding to spurious patterns within data, and limited transparency (Lones, 2021), not to mention higher training and inference times — so in general are not likely to be a productive approach for malware detection unless they lead to a significant improvement in performance.
Table 31. Overall best-performing models
ID Feature Set Type Model Feature Selection Accuracy F1-Score Precision TPR TNR
$C_{1}$ API Static Random Forest Mutual information 0.967±$\pm 0.001$ 0.969±$\pm 0.001$ 0.967±$\pm 0.002$ 0.963±$\pm 0.002$ 0.970±$\pm 0.002$
$C_{2}$ API Static Random Forest Variance threshold 0.968±$\pm 0.002$ 0.971±$\pm 0.002$ 0.968±$\pm 0.002$ 0.965±$\pm 0.002$ 0.971±$\pm 0.001$
$C_{3}$ API Static Random Forest Chi-square 0.961±$\pm 0.003$ 0.915±$\pm 0.004$ 0.954±$\pm 0.002$ 0.967±$\pm 0.001$ 0.953±$\pm 0.001$
$C_{4}$ Permissions Static Random Forest Mutual information 0.930±$\pm 0.001$ 0.932±$\pm 0.001$ 0.929±$\pm 0.002$ 0.931±$\pm 0.002$ 0.931±$\pm 0.003$
$C_{5}$ Permissions Static Random Forest Variance threshold 0.949±$\pm 0.003$ 0.948±$\pm 0.002$ 0.950±$\pm 0.001$ 0.951±$\pm 0.001$ 0.949±$\pm 0.002$
$C_{6}$ Opcodes (5-gram) Static Random Forest Mutual information 0.973±$\pm 0.002$ 0.979±$\pm 0.002$ 0.975±$\pm 0.001$ 0.965±$\pm 0.001$ 0.979±$\pm 0.002$
$C_{7}$ System Calls (3-gram) Dynamic Random Forest Chi-square 0.951±$\pm 0.002$ 0.951±$\pm 0.001$ 0.950±$\pm 0.002$ 0.949±$\pm 0.003$ 0.952±$\pm 0.001$
$C_{8}$ System Calls + API Calls (Usage) Dynamic Random Forest None 0.961±$\pm 0.002$ 0.964±$\pm 0.002$ 0.963±$\pm 0.001$ 0.957±$\pm 0.003$ 0.965±$\pm 0.002$
$C_{9}$ Network Traffic (TCP) Dynamic Random Forest None 0.977±$\pm 0.003$ 0.969±$\pm 0.003$ 0.981±$\pm 0.003$ 0.978±$\pm 0.002$ 0.971±$\pm 0.002$
$C_{10}$ Network Traffic (TCP + HTTP) Dynamic kNN None 0.967±$\pm 0.004$ 0.964±$\pm 0.002$ 0.967±$\pm 0.001$ 0.969±$\pm 0.002$ 0.964±$\pm 0.003$
$C_{11}$ Network Traffic (TCP) + Permissions Hybrid Random Forest None 0.943±$\pm 0.001$ 0.941±$\pm 0.002$ 0.945±$\pm 0.002$ 0.948±$\pm 0.002$ 0.940±$\pm 0.002$
$C_{12}$ Permissions + System Calls Hybrid Random Forest Mutual information 0.948±$\pm 0.02$ 0.948±$\pm 0.03$ 0.949±$\pm 0.03$ 0.935±$\pm 0.01$ 0.961±$\pm 0.02$
7. Ensemble Models
In the hybrid approaches discussed in the previous section, a single ML model was built from a feature set containing more than one type of feature. However, a limitation of this approach is that all the features must be used with the same model, and must be represented in a way that is compatible with the same model. In this section, we explore a more flexible approach: ensemble modelling. Ensembles are ML models that combine the outputs of other, potentially diverse, ML models. A general benefit of ensembles is that they are often able to combine the strengths of their component models in order to provide a more robust overall prediction.
In this section, we explore whether better models can be created by combining the best standalone malware classification models. In particular, we construct voting ensembles from the best-performing classifiers listed in Table 31. Voting ensembles are constructed by selecting an odd number of these so-called base models, and then outputting the majority decision, so they do not require further training. We considered each odd combination of base classifiers, and then sorted them by accuracy. Table 32 lists the top five ensemble models according to this criterion.
Table 32. Performance of ensemble models, showing the top five voting ensemble models formed from the overall best-performing base classifier models ($C_{n}$) listed in Table 31.
ID Classifiers Accuracy F1-Score Precision TPR TNR
$E_{1}$ $C_{1}$ + $C_{3}$ + $C_{7}$ 0.978±$\pm 0.02$ 0.976±$\pm 0.03$ 0.975±$\pm 0.01$ 0.979±$\pm 0.01$ 0.977±$\pm 0.03$
$E_{2}$ $C_{2}$ + $C_{3}$ + $C_{7}$ 0.971±$\pm 0.02$ 0.971±$\pm 0.02$ 0.970±$\pm 0.02$ 0.971±$\pm 0.02$ 0.967±$\pm 0.02$
$E_{3}$ $C_{1}$ + $C_{4}$ + $C_{6}$ 0.959±$\pm 0.03$ 0.955±$\pm 0.03$ 0.953±$\pm 0.02$ 0.956±$\pm 0.03$ 0.952±$\pm 0.04$
$E_{4}$ $C_{3}$ + $C_{4}$ + $C_{9}$ + $C_{10}$ + $C_{12}$ 0.969±$\pm 0.01$ 0.970±$\pm 0.04$ 0.965±$\pm 0.04$ 0.966±$\pm 0.01$ 0.969±$\pm 0.03$
$E_{5}$ $C_{4}$ + $C_{5}$ + $C_{8}$ + $C_{9}$ + $C_{10}$ 0.954±$\pm 0.02$ 0.953±$\pm 0.01$ 0.952±$\pm 0.02$ 0.955±$\pm 0.01$ 0.958±$\pm 0.03$
Whilst all of these models perform better than the hybrid models presented in the previous section, Kruskal-Wallis and Dunn’s tests showed that none of them performed significantly better than the best standalone network traffic model. However, the $E_{1}$ ensemble does perform better than the best API call model, and, importantly, it achieves this by combining non-network features. This is significant because it avoids the challenges (discussed in Section 4.3) of recording network traffic during the dynamic analysis process, resulting in a model that may be more practical for real world deployment.
8. Discussion
This study allowed us to take a broader perspective on the Android malware-detection literature and, arguably for the first time, see how the different ML-based approaches to solving this problem work within contemporary Android systems. Some of the observations were surprising. For instance, our results suggest that there is generally little advantage to using dynamic analysis techniques, and that very accurate models can be built using relatively simple static features, particularly API call usage. The only scenario where dynamic analysis seemed to bring an advantage was when we built models based around network activity. However, this is expensive to implement in practice, since collection of network activity requires dynamic analysis at both training and inference time, and relies on features which are both difficult to collect and potentially brittle.
This is not the only instance where we observed a trade-off between the difficulty of extracting features, and the accuracy of the resulting models. Another notable example is the use of API calls versus permissions. Generating features based upon API calls requires the use of a static analysis tool. By comparison, permissions can be read directly from an application’s manifest file, a trivial operation. However, we found models trained on API call features to be significantly more accurate than models based upon permissions, showing that it is sometimes worth expending more effort on feature extraction.
We also found that hybrid models, built from a combination of static and dynamic features, were no better than those built from static or dynamic features alone. However, importantly, we showed that more accurate models can be constructed by ensembling the best static and dynamic models. Most significantly, we found that ensembling can be used to build models with accuracies comparable to the best network traffic models, but using less brittle features.
A recurring observation was the challenge caused by the growing complexity of Android. For instance, we observed that the number of API calls has grown by two orders of magnitude since early work on ML-based Android malware detection. This means the use of feature selection is now very important in order to reduce feature numbers down to manageable levels. Within this context, an important observation is that feature selection does not just reduce the computational resources needed for training; it also, in most cases, increases the accuracy of the resulting models. Most notably, we have shown that reducing the number of API calls by 95% leads to higher levels of malware discrimination. In general, we found mutual information and chi-square to be the most effective feature selectors, and did not observe any benefit to using a more complex model-based feature selection approach. The Drebin feature set has historically played an important role in ML-based malware detection studies. Although this feature set still performs reasonably well, our results suggest that it is no longer the best feature set to use, and is difficult to deploy due to the large number of features it entails when used within a contemporary Android environment. Unlike other feature sets we studied, it does not benefit from feature selection.
For features such as API calls, where it is possible to measure frequency of use in addition to whether they are used, this extra information generally does not improve a model’s ability to discriminate malware. This is a somewhat unexpected finding, but was observed consistently across a number of feature sets, both static and dynamic.
Beyond features, we also found that model choice is important, and random forests were surprisingly consistent in being the best performing models across the majority of experiments and feature sets. SVMs also performed well in general, but were more expensive to train than random forests. Naïve Bayes models generally performed poorly, suggesting that, for malware detection, interactions between features are important. A somewhat surprising observation, given their increasing prevalence within ML more generally, is that deep neural network models did not perform particularly well. However, this may be due to the use of pre-extracted features in the majority of malware detection studies. In situations where raw data is used (e.g. raw source code), there may be benefits to using more complex deep learning models such as transformers. However, there are also significant risks of overfitting and greater sensitivity to distribution shifts (Lones, 2021).
Finally, we observed that a number of the malware detection approaches reported in the literature performed significantly less well when we reimplemented and evaluated them on a large contemporary dataset. In many cases, this may be attributed to the increasing complexity of the Android environment; however, poor experimental practice is also likely to be a contributing factor, and we have observed many instances of authors falling into common ML pitfalls (Lones, 2021), including data leakage and the use of inappropriate metrics. Either way, it shows the importance of periodically reimplementing and reevaluating studies.
9. Conclusions
In this work, we have have experimentally investigated the factors that affect model accuracy when applying machine learning to Android malware detection. We focused, in particular, on the importance of feature and model selection. An important component of our investigation was the reimplementation and reevaluation of existing approaches from the literature. This allowed us both to carry out a rigorous experimental comparison of these approaches, and to determine whether they remain relevant to contemporary Android systems. We also carried out a number of new experiments to fill in the knowledge gaps found in previous studies. These experiments were supported by the collection of a large up-to-date data set comprising 124,000 Android applications, equally divided between malware and benign applications. The tools used to create this dataset, as well as the tools we developed to extract features, are provided as a resource for the community.
A key observation is that features extracted using static analysis of Android APKs are often more effective than those extracted using dynamic analysis of running applications. This is important to know, since static analysis is much cheaper than dynamic analysis. Within this context, we found that simple representations of API usage led to very effective models, with accuracies of up to 96.8%. The exception to this is when features extracted from network traffic are used to build models. In our study, it is these models that provided the highest levels of malware discrimination, achieving accuracies of up to 97.7%. However, they are expensive to train and use, and rely on features that are difficult to collect and potentially brittle. Fortunately, we also showed that ensemble models built from non-network features can achieve a similar level of discrimination (97.8%), offering a more practical alternative for real world deployment.
Other notable observations include the importance of using feature selection algorithms when dealing with the complexity of modern Android systems, the high accuracy of random forest models in comparison to other ML models (including deep learners), and the relatively poor performance of the historically-important Drebin feature set. In future work, we plan to focus on online ML-based Android malware detection. Keeping a model up to date is important for accurate malware detection, and online learning can also be useful for drift detection. The sheer number of applications released everyday makes Android a prime candidate for online machine learning.